A Bibliography of Publications of Friedrich Ludwig Bauer

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 February 2024 Version 1.10

Abstract This bibliography records publications of Friedrich Ludwig Bauer (10

June 1924–26 March 2015).

Title word cross-reference

$24.50 [G.70]. 3, ({2p−1 \atop p−1}) ≡ 1 (mod p^{3}) [Bau88a]. e^{π} [Bau07c, Bau07c]. g [Bau60b]. π [Bau05a]. π(x) [Bau03d]. π^{e} [Bau07c]. QR [RB68].

-algorithm [Bau60b].

1 [Bau05c, Bau06d]. 105 [BS62c]. 10th [BS81]. 11 [BGHS79]. 11/27/1978 [BGHS79]. 135 [Sch69]. 153 [Bau63a]. 17-Eck [BH08]. 17-gon [BH08]. 1905 [Bau05c, Bau06d]. 1945 [Bau08b]. 1958 [Lan59]. 1960 [AGG60]. 1964 [Ghi10]. 1971 [Dij71, FGR72]. 1972 [BW72a]. 1973 [Bau76b]. 1978 [BGHS79]. 1980 [GB81]. 1988 [Bau89b]. 1989 [Bau91a]. 1990 [Bau90c]. 1991 [BBS93]. 1999 [BS00].

2001 [Bau01a]. 2007 [Bau07e]. 20th [BS00]. 22.8.1970 [GB71]. 25.5.1980 [GB81]. 28/29 [BD01]. 292 [Chu51].

3 [Bau95a, Leh87]. 3-211-82527-4 [Bau95a]. 3-540-13814-5 [Leh87]. 300th [Bau07c, Bau07g]. 310447Theoretische [Bau95a]. 314 [Čul72b]. 367 [Čul72a].

4 [Bau95a]. 4/21/1909 [BGHS79].

5 [Leh87]. 50 [Gep66].

60 [BBG^{+}60, BBG^{+}63c, BBG^{+}63a, BBG^{+}76, Bau61a, BHO^{+}67, Čul72b, Čul72a, Gep66, GHL67, NBB^{+}60, NBB^{+}62, NBB^{+}63, NBB^{+}65, NBB^{+}97]. 68 [vWMPK69b, vWMPK69a].

80 [Ano04, Bod04, Hei68, Kad76].

Abb [Bau65a, Hei68, Leh87]. abbreviated [Bau56d]. abbreviation [Bau55a]. abgekürzten [Bau56d]. Abhandlungen [BBG^{+}58]. Abkürzung [Bau55a]. Abrißder [Bau90c]. Abschätzung [Bau65e]. Abschätzungen [BDS69]. Absolute [BSW61, BH61]. abstract [Bau85b, PBB^{+}82]. Abstracts [Bau74d]. abstrakte [Bau85b, PBB^{+}82]. Academy [Bau00, Bau00f]. accuracy [Bau66a, Bau06a, SB53]. achieve [SB54]. Ackermann [BBG^{+}58]. Acta [BB03]. Advanced [Bau73a, BDG^{+}74, BDW^{+}75, Bau77c, BBS93, Scr79, BE76, Ghi74]. Aerodynamik [MBH^{+}56]. Aerodynamik/Boyd [MBH^{+}56]. After [Bau09c, Bau10c, Bau10b]. aided [BMPP89]. Aiken [Bau98c, Bau98c]. aims [BS60a]. AISMC [CCP96]. AISMC-3 [CCP96]. Akademie [Bau65a, Bau00, Bau00f, Gep66]. Akademie-Verlag [Bau65a, Gep66]. Aktuelle [Bau57b]. ALCOR [SB62]. Alcune [Ghi10]. Algebra [Bau57b, Bau74d, WR71, Bau49b, Bau54c, Bau81, Bau82a, Bau88b, Bau91a, BBS93, Bau88b]. algebraic [Bau54f, Bau56a, Bau56d, Bau57c, BW88, PBB^{+}82]. Algebraik [Bau49b, Chu51]. algebraische [Bau56d, PBB^{+}82]. algebraischen [Bau54f, Bau56a]. algebraischer [Bau57c]. algébre [Bau54c]. ALGOL [BBG^{+}63c, Čul72b, Čul72a, HB62, NBB^{+}62, Rut13, Gep66, BBG^{+}63c, BBG^{+}63a, BBG^{+}63b, BBG^{+}76, BS60a, BHO^{+}67, Bau11a, Bau12, GHL67, HB62, NBB^{+}60, NBB^{+}62, NBB^{+}63, NBB^{+}65, Rut13, vWMPK69b, vWMPK69a, BBG^{+}60, BFBS64, NBB^{+}97, Car66, Hoh64, Rhe65]. ALGOL-60 [BBG^{+}76, NBB^{+}65]. ALGOL-Verschwörung [Bau11a, Bau12]. ALGOL-Wörterbuch [HB62]. algorithm [Bau60b, Bau61a, BS62c, Bau63a]. Algorithmic [BBG^{+}63c, BBG^{+}76, BW82, NBB^{+}60, NBB^{+}63, NBB^{+}97, vWMPK69b, vWMPK69a, BBG^{+}60, BBG^{+}63a, BBG^{+}63b, Bau97d, Čul72b, Čul72a, Gep66, NBB^{+}65, BW81]. Algorithmische [BBG^{+}63c, Gep66, NBB^{+}62, BW81]. Algorithms [Bau65f, Bau81, Bau82a, Bau10c, Bau10a, EK81, Bau59c]. Allgäu [Bau03a, Bau03a]. allgemeinen [Bau55a]. Allies [Bau00d]. Alliierten [Bau00d]. Allocation [Bau90a]. Alps [Rut13]. also [Čul72b, Čul72a]. am [Bau74d]. American [BBG^{+}58]. analisi [Ghi10]. analysis [Ghi10]. Andrei [Bau89b]. Anfängen [Bau08d]. Angelitch [G.70]. angewandten [Knö66]. Angriff [Bau00d]. Angstl [Bau77a]. Angular [MBH^{+}56]. Anläßlich [Bau07c, Bau07g, BS81]. anniversary [BS81]. Anschauliche [BBG^{+}58]. Antimon [BHB50]. Antimon/Streuversuche [BHB50]. antimony [BHB50]. antimony/scattering [BHB50]. Anwendung [BHS^{+}56]. Anwendungsgebiete [Sch70, Sch69]. apparatus [Bau59b]. Appendix [Bau74a, Bau77b]. Approach [BB84, Bau79b, Ion84, PPWB78]. approximate [Šve70]. approximation [Lan59]. approximations [Bau06a, Bau07a]. April [Bau89b, Lan59]. Arbeit [Rut13]. Arbeitsblatt [DS76]. Arbeitstagung [KB53, KB53]. Arbib [Ion84]. Arccotangent [Bau07g]. Archduke [Bau07f]. architecture [Roj96]. Architectures [RH00]. Architektur [Roj96]. Arcus [Bau07g]. Arcus-Cotangens-Relationen [Bau07g]. arising [Šve70]. arithmetic [PBB^{+}82]. articles [Hof62]. Artificial [CCP96]. artige [Bau08e]. aspects [BRS63, Bau68b, Rut69]. Association [Čul72b]. Atmosphere [MBH^{+}56]. Atmosphere [?]DerMoritz:1956:BBM. Atombau [MBH^{+}56]. Atombau/Schulz [MBH^{+}56]. Atombau/Schulz-Methke [MBH^{+}56]. atomic [Bau51]. Atomkerne [Bau51]. Atomphysik [BHS^{+}56]. Atomphysik/Sears [BHS^{+}56]. attack [LMNS16]. attacking [Bau00d]. Aufgaben [BGH75, BGH76]. Auflage [HB62]. August [BS76, Bau91a, BBS93, BS00, Bau05c, Bau06d, FGR72, Bau05c, Bau06d]. Aumann [G.70]. ausgewählten [Hof62]. Ausreduzierte [Bau49a]. Aussagenlogik [BW91]. Ausschusses [KB53]. Austria [CCP96]. Ausübung [BS57a, BS60b, Bau01b]. Automatic [BS62a, Sch69, Sch70, BS57a, BS60b, Bau01b]. automatischen [BS57a, BS60b, Bau01b]. award [Bau11c].

B [Čul72b, Čul72a, Gep66, Ghe87, Sch70]. Babbage [SBET85]. Backus [Čul72b, Čul72a, Gep66]. Band [Bau65a, Dij71, Knö66, Sch69, Sch70]. Bau [BHS^{+}56]. Bauer [Ano04, Bod04, Car66, Chu51, Čul72b, Čul72a, Dij86, G.70, Gep66, Ghe87, Hei68, Hoh64, Ion84, Kad76, Knö66, Leh87, Liv81, Pet99, Rhe65, Sch69, Sch70, Scr79, Ale06, Ano05, Ano16, Ano19, AB87, BKB13, Bro15, Chr07, wEC86, wC03, wCL04, DZ75, Dij71, Dij72, DS76, FGH^{+}06, Gan15, Ghi74, Kil12, Kob10, LMNS16, Li88, MB96, Mol15, NS64, OML09, Pet99, Pry92, Rut69, SW12, SBE^{+}99a, Šve70, Zen68]. Bauer/G [Kad76]. Baum [Bau08a]. Baumann [Car66, Hoh64, Rhe65]. Bavarian [Bau00, Bau00f]. Bayerischen [Bau00, Bau00f]. bearbeitet [Hei68]. Beauty [FvGGM90]. Bechert [BHS^{+}56]. Beginnings [Bau08d]. Begins [Bau09c]. Behandlung [Bau56a]. Behavior [BHS^{+}56]. Beiträge [BS81, Bau54b, Bau56c, Bau56b, Bau57a]. Beiträgen [Hof62]. Bekker [BBG^{+}58]. Bemerkung [Bau89a]. Berechnung [Bau65e]. Berghammer [Ghe87]. Bericht [BBG^{+}63c, Gep66, NBB^{+}62]. Berichtigungen [BG51]. Berlin [Bau65a, G.70, Gep66, Kad76, KB53, Leh87]. Berlin-Heidelberg-New [Kad76, Leh87]. Berlin/Theoretical [KB53]. Berlin/Theoretical-physical [KB53]. Berlin/Theoretisch [KB53]. Berlin/Theoretisch-physikalische [KB53]. Bernoulli [Bau54b, Bau54b, Bau54a, Bau58a]. Bernoulli-Jacobischen [Bau54b]. Bernoullischen [Bau54a]. Berücksichtigung [Ale06, Sch69, Sch70]. besiegte [Bau05c, Bau06d]. besonderer [Sch69, Sch70]. Bestimmung [BBG^{+}58]. Betriebs [Bau55b]. Betriebs-Kettenmatrix [Bau55b]. Between [Bau80, Bau57b]. Betz [MBH^{+}56]. Bildung [BHS^{+}56]. Binary [Gri90]. Biografie [Rut13]. Biographies [Wei88]. biography [Ano19, Rut13]. birth [Dij88]. Birthday [FvGGM90, Bau07c, Bau07g]. Blaschke [BBG^{+}58]. Blindenschrift [Bau96c]. Bombelli [Bau06e, Bau06e]. bombes [Bau07b]. Bonn [BD01]. Book [Ale06, Bau95a, Car66, Čul72b, Čul72a, Dij72, G.70, Ghe87, Ghi74, Hoh64, Ion84, Knö66, Kob10, Leh87, Liv81, Rhe65, Sch69, Sch70, Scr79]. born [Bau05c, Bau06d, GB71, GB81]. Boyd [MBH^{+}56]. Braille [Bau96c]. Braun [GTB53]. breakthroughs [Bau04c]. Brief [Bau90c]. British [SBET85]. Brockhaus [MBH^{+}56]. Broglie [Bau54c]. Broy [Ghe87, Liv81]. Brückner [Ghe87]. Bruhat [BBG^{+}58]. Buchbesprechung [Bau65a, Gep66, Hei68, Kad76]. Bücher [BHS^{+}56, BBG^{+}58, GTB53, MBH^{+}56, BG51, BHS^{+}56, BBG^{+}58, GTB53, MBH^{+}56]. Bulirsch [Ale06, G.70]. Bulletin [BBG^{+}63c, BBG^{+}63c]. Business [FvGGM90].

C [Bau95a, Čul72b, Čul72a, Gep66, Ion84]. C.I.M.E [Ghi10]. calculating [BS57a, BS60b, Bau01b, Roj96]. calculation [Bau65e, SB53]. Calculations [Bau10d, Bau10f]. calculator [Bau98b]. calculators [SB53, SB54]. calculus [Bau49b]. Call [Ano05]. Cap [Bau95a]. Capsule [SBET85]. Carl [BH08]. carry [BS57a, BS60b, Bau01b]. carryover [Bau98c]. Cataldi [Bau06e, Bau06e]. Cellar [Bau90a]. Central [Bau80]. Centro [Ghi10]. Century [MHR80]. Certain [HB59, HB60, BF58, Lev89]. chain [Bau55b, Bau08e]. Change [Bau09f]. Characteristic [BH60a, HB59, BH61]. Chast′ [BG90]. checking [Bau77a]. Chemie [BBG^{+}58]. Chemie/Spitzer [BBG^{+}58]. Chemischer [BBG^{+}58]. choice [Bau61c]. choix [Bau61c]. Choquet [Che74]. CIP [Ghe87, Bau79c, BBB^{+}85, BEH^{+}87]. CIP-S [BEH^{+}87]. cipher [LMNS16]. Claude [Bau01a]. clear [SB54]. Cleve [Mol15]. cm [G.70]. coded [BS57a, BS60b, Bau01b]. Codes [Bau07i, SBET85, Bau07i]. collaboration [BW81]. Collatz [Bau08a]. Collatz-Baum [Bau08a]. Collection [BBD^{+}81, MHR80, BGH81]. College [Bau76b]. Colloquium [BS81]. Combinations [Bau65c]. comma [Bau96d]. Commemoration [Bau07e]. Comments [DS76, Bau64]. Commission [Bau00, Bau00f]. Committee [KB53]. common [BS60a]. Communications [Čul72b]. Commuting [Pry92]. Compiler [Bau74a, BDG^{+}74, BE76, Bau77b]. Comprehensive [dLB07]. computability [Ion84]. Computation [Sch70, Bau91a, BS00, CCP96, Sch69]. Computational [Bau74b, Rut69]. Computer [Bau60a, Bau00, Bau07j, Bau08d, Čul72a, Gan15, Ion84, Leh87, Liv81, Bau74c, BMPP89, Bau99c, Bau06c, Bau09d, EK81, SBE^{+}99b, BG73, BG74, BGH75, BGH76, BG82, BG84, BG91, BG92, Bau06c]. Computers [RH00, Bau87, HB62]. Computing [BBR^{+}58, BS62a, Bau80, Bau10h, Čul72b, MHR80, Bau54b, Bau56c, Bau57b, BHSS65, Lev89, Kil12]. Conclusion [Bau10e]. Concurrent [BB84]. Condition [Bau63b, BH60b, Šve70]. Conference [BS60a, BD01, CCP96, KB53]. conformal [Šve70]. Congress [FGR72]. Conjecture [Dij86]. connection [OML09]. Conspiracy [Bau11a, Bau12]. constant [Bau89a]. constitution [Bau11c]. Construction [Bau74a, BDG^{+}74, Bau77b, BB79, BE76, BMPP89, Liv81]. Constructs [BB84]. Continued [Bau65f, Bau06e, Bau09e, Bau10a, BKB13, BF58, Bau03d, Bau05b, Bau06a, Bau09e, Bau09f]. Contributions [BD01, Bau54b, Bau56c, Bau56b, Bau57a, BS81, BD01]. Controlled [Bau60a, Bau54b, Bau56c, EPBS63]. convergence [Bau54a, RB55]. convergent [Bau54b, Bau54f, Bau55a]. converse [Li88]. convexity [Zen68]. Cooperation [Bau10h]. Corner [Mol15]. Correcting [Bau07i, SBET85]. Correspondence [BH59]. Cosine [Bau05a]. Cosinus [Bau05a]. Cotangens [Bau07g]. Courant [BBG^{+}58]. Course [Bau73a, BDG^{+}74, BDW^{+}75, Bau77c, BE76, Bau76b, Ghi74, Scr79]. Crypt [BW88]. Crypt-equivalent [BW88]. Cryptological [Bau96a]. Cryptology [Bau97b, Bau00a, Bau00b, Bau00c, Bau02a, Bau07d, Chr07, Bau95c, Bau97c, Bau82b, Bau93, Bau94b, Bau96c, Kob10]. cubes [Bau04b]. Current [Bau57b]. Curriculum [FGH^{+}06]. cybernetics [LB69]. Czech [MBH^{+}56].

Danilewski [Bau57a]. Danilewski-Verfahren [Bau57a]. dans [Bau54c]. Darmstadt [Hei68]. d’arrondi [Bau61c]. data [BS57a, BS60b, Bau85b, Bau01b, PBB^{+}82]. Daten [BS57a, BS60b, Bau01b]. Datentypen [Bau85b, PBB^{+}82]. David [Ion84]. Davis [BHS^{+}56]. Dead [Bau04c]. death [BS81]. December [Bau74d, Bau89b]. Decimal [Gri90]. Deciphered [Bau97c, Bau09a]. deciphering [Bau08b, Bau95c, Bau00c]. decomposition [Bau55c]. Decrypted [Bau97b, Bau00a, Bau00b, Bau02a, Bau07d, Chr07, Kob10]. dedicated [Bro15]. defeated [Bau05c, Bau06d]. Definite [BR71]. Dependent [Bau10e]. Description [BBR^{+}58, BHO^{+}67]. Design [Bau78]. Detailization [Bau79a]. Détermination [RB55, RB55]. determining [Bau54b, Bau56d]. Deutsche [BD64]. Deutsches [Bau96a]. Development [BBG^{+}78a, Bau80, BW82, Bau54b, Bau56c, BBG^{+}78b, Bau79c, BBG^{+}79, BW81]. Devices [Bau96a, Bau99b, Bau10b]. dictionary [HB62]. died [GB71, GB81]. Dielectric [BHS^{+}56]. difference [Bau59c]. different [BB49]. Digital [Bau80, HB62, Hof62]. Digital-Rechenanlagen [HB62]. Digitale [Hof62]. Dijkstra [BB03, FvGGM90]. d’intégration [Bau61d]. Dirac [Bau52c]. Direct [Bau56b, Bau55c, Bau56c]. Direkte [Bau56c, Bau56b]. direktes [Bau55c]. discs [DZ75]. disks [Bau68a]. disturbances [Bau53]. divisors [Wil84]. DM [Hei68, Kad76, Leh87]. doctorates [Bau11c]. Documentation [KB53]. does [Bau74c]. Dokumentation [KB53]. dont [Bau52c]. Dosch [Ghe87]. double [Bau07a]. Dr [DS76]. dry [OML09]. dual [Bau94a]. d’une [RB55]. dünner [BHS^{+}56]. durch [Bau92, HB62]. Durchbrüche [Bau04c]. Durchführung [Bau54b].

Early [Bau80, Bau07j, Bau06c, Bau06b]. Eck [BH08]. Edited [BBG^{+}63c]. Edition [Chr07, Kob10, HB62]. editor [Bau95b, Čul72b, LYB98, RBHM95]. Edsger [BB03, FvGGM90]. Eduard [BGHS79]. Efficient [Bau90b]. eigen [Bau61e]. eigenvalue [Bau54f, Bau56a, Bau56d, Bau57c]. Eigenvalues [wC03, Bau65e, BDS69, wCL04]. eigenvectors [RB55]. Eigenwerte [BDS69]. Eigenwerten [Bau65e]. Eigenwertprobleme [Bau56d, Bau57c]. Eigenwertproblemen [Bau56a, Bau54f]. einführende [BG71a, BG71b, BG73, BG74, BG84, BG91, BG92, Kad76]. Einführung [Ale06, BHSS65, BBG^{+}58]. einigen [Bau57b]. einschrittige [Bau98c]. Einzeldarstellungen [Sch69, Sch70]. Electromechanical [Bau10b, Bau06c]. Electronic [Bau10c]. Elektroakustik [MBH^{+}56]. Elektroakustik/Betz [MBH^{+}56]. elektromechanischer [Bau06c]. Elektronenstrahl [MBH^{+}56]. Elektronenstrahl-Oszillograf [MBH^{+}56]. Elektronenstrahl-Oszillograf/Reichardt [MBH^{+}56]. Elektronik [BHS^{+}56]. Elementare [BW91]. Elementarteilchen [Bau49a, BHB50, BB49]. Elementarteilchen/Kernsplitterung [BHB50]. elementary [Bau49a, BHB50, Bau65b, Bau06a, BB49, Wil84, BW91]. elements [Bau52c, Bau52c]. Elimination [Bau65c, Bau08c, Bau08c]. Elwood [Bau01a]. Encryption [Bau99b]. ends [Bau04c]. energetic [BHB50]. energiereichsten [BHB50]. Engineer [AAB^{+}68]. Engineering [Bau73a, Bau73b, BDW^{+}75, Bau77c, BD01, Scr79, Bau72, Ghi74]. ENIGMA [Bau05c, Bau06d, Bau05c, Bau06d, Bau00d]. Entwicklung [Bau54b, Bau56c]. Entzifferte [Bau00c, Bau95c, Bau97c, Bau09a]. Entzifferung [Bau08b]. Episoden [Bau08d]. Episodes [Bau08d]. epsilon [Bau59c]. Equations [Bau65c, Bau49a, Bau52d, Bau52a, Bau54f, Bau66a, BB49]. equilibrium [Che72]. equivalent [BW88]. Erckmann [BBG^{+}58]. erfand [Bau98b]. Ergänzungen [BG51]. Erich [Bau08b]. Erinnerung [BBG^{+}63c]. Erinnerungen [BS81]. erreur [Bau61c]. Error [Bau74b, Bau99b, Bau07i, SBET85, Bau61c]. Error-Correcting [SBET85]. Ershov [Bau89b]. Erster [BG73, BGH75, BG82, Dij71]. erweiterte [HB62]. Erzherzog [Bau07f]. Erzielung [SB54]. especially [BS60a]. Essays [MHR80]. Estimates [BDS69]. estimation [Bau65e]. Estivo [Ghi10]. ETH [Bru11, Bau08d, Bru11]. ETH“ [Bau08d]. Euclidean [BH60b]. Euler [Bau07c, Bau07g, Bau89a, Bau07c, Bau07g]. Eulersche [Bau89a]. Europe [Bau80]. evaluation [Bau79a]. evolutionary [Bau76c, Bau76d]. Examples [BBD^{+}81]. Exclusion [BF60]. execution [Bau54b]. exercises [BGH75, BGH76]. Expanding [HB59]. experiments [BHB50]. Exploration [MBH^{+}56]. extended [HB62].

F [Ale06, Car66, Chr07, Čul72b, Čul72a, DS76, G.70, Gep66, Ghe87, Ghi74, Hei68, Hoh64, Ion84, Kad76, Leh87, Liv81, Rhe65, Sch69, Sch70]. F. [Ano04, Bod04, Chu51, Dij86, Knö66, Pet99, Rut69, Sch69, Sch70, Scr79, SBET85]. Fachbegriffe [HB62]. factorial [Bau07a]. Factoring [Bau56b]. factorization [Bau56c]. facts [Bau96b]. Fakten [Bau96b]. Faktorisierung [Bau56c, Bau56b]. fate [Bau09b]. fates [Bau97e]. February [AGG60]. Feenberg [MBH^{+}56]. Fehlerkorrigierende [Bau07i]. Feldherr [Bau04e]. Feldmechanische [BB49]. Feliciano [Car66, Hoh64, Rhe65]. Feshbach [MBH^{+}56]. Festschrift [Bau07e]. field [Bau62, Zen68, BB49]. Fields [Bau68a, NS64, BHS^{+}56]. Fields/Kaplan [BHS^{+}56]. Fike [wEC86, wC03, wCL04, Li88, Pry92, SW12]. First [RH00, Bau55a, BG73, BG82]. five [HB62]. Fleissner [Bau07f]. Fleissner-Raster [Bau07f]. floating [SB53]. forerunner [BW72b, BW72a]. Form [Bau59a]. Formal [BMPP89, Bau10c, BF58, Bau79b, EPBS63]. formaler [Bau66c]. Formelübersetzung [SB59, SB01]. Formula [Bau60a, SB60, SB83, Bau07a, SB59, SB01]. Formula-Controlled [Bau60a]. formulae [Bau77a]. Formulas [Bau97a, Bau66c]. Foundations [BS00, Bau10h, Kil12]. Fourier [BS59]. Fourier-Transformation [BS59]. Fourth [Chr07, Kob10]. Fraction [Bau09e, Bau09f, Bau03d, Bau05b]. Fractions [Bau65f, Bau06e, Bau09e, Bau10a, BKB13, BF58, Bau06a]. fragmentation [BHB50]. Franksen [SBET85]. free [Bau77a]. French [Bau52b, Bau52c, Bau54c, Bau61c, Bau61d, RB55]. Frequenz [BBG^{+}58]. FRG [Bau91a]. Friedrich [Bro15, Kil12, Kob10, Ano05, Ano16, Ano19, AB87, FGH^{+}06, Gan15, Mol15, SBE^{+}99a]. Friedrichs [BHS^{+}56]. Fritz [Bau09b]. Fromme [Bau04d]. Frössel [BG51]. Frühe [Bau06b]. früher [Bau06c, Bau07j]. Fu [Ion84]. Fully [BBG^{+}58]. fünf [HB62]. Funktionen [GTB53]. Funktionen/Sommerfeld [GTB53]. Further [LMNS16, Bau61b]. fusion [Bau54c].

G [G.70, Kad76]. Gases [BBG^{+}58]. Gases/Wirk [BBG^{+}58]. Gatlinburg [Bau74d, MB96]. Gauß [BH08, BR71, BH08]. geb [Bau65a, GB71, GB81, GB81]. gebrauchte [Bau05c, Bau06d]. Geburtstags [Bau07c, Bau07g]. Gedächtniskolloquium [BS81]. Gedenken [Bau05c, Bau06d]. gegen [Bau00d]. Geheimnisse [Bau95c, Bau97c, Bau00c, Bau09a]. Geiselbrechtinger [Ghe87]. Geleitworten [Leh87]. Genauigkeitsfragen [Bau66a]. general [Bau55a, Bau04e]. Generalization [wEC86, Bau61b]. generalized [DZ75]. generator [EPBS63]. Genesis [Bau97d]. Geometrie [BBG^{+}58]. Geometrie/Ackermann [BBG^{+}58]. German [AAB^{+}68, BBG^{+}63c, Bau49b, Bau49a, Bau50, BHB50, Bau51, Bau52d, Bau52a, Bau54b, Bau54f, Bau54a, Bau54e, Bau55b, Bau55c, Bau55a, Bau56c, Bau56b, Bau56a, Bau56d, Bau57a, BS57b, BS57a, Bau57c, Bau57b, BS59, BS60b, Bau61e, BS62b, BD64, BHSS65, Bau65e, Bau66a, Bau66c, Bau68c, BDS69, BW72a, BG73, BG74, Bau74c, BGH75, BGH76, Bau76e, BW81, BS81, BG82, Bau82b, BG84, Bau85a, Bau85b, Bau88b, Bau89a, Bau90c, BW91, BG91, Bau91b, BG92, Bau92, Bau93, Bau94b, Bau95c, Bau96c, Bau96b, Bau96d, Bau96e, Bau97c, Bau97e, Bau00, Bau98b, Bau98c, Bau99c, Bau99a, Bau00c, Bau00f, Bau00d, Bau01a, Bau01b, Bau03a, Bau03c, Bau04a, Bau04b, Bau04c, Bau04d, Bau04e, Bau05b, Bau05c, Bau05a, Bau05e, Bau06b, Bau06d, Bau06e, Bau06c, Bau07c, Bau07e, Bau07f, Bau07g, Bau07i, Bau07j, BH08, Bau08d, Bau08b, Bau08c, Bau08a, Bau08e, Bau09a, Bau09b]. German [Bau09d, Bau09e, Bau09f, Bau11a, Bau11c, Bau12, BB49, Bru11, Dij88, GB71, GB81, HB62, Hof62, KB53, LB69, NBB^{+}62, PBB^{+}82, Pet99, Roj96, Rut13, SB53, SB54, SB59, SB01, SBE^{+}99b, SB90, Sto93, Sto94]. Germany [Bau74d, BBS93, BS00]. Gershgorin [Bau68a, DZ75]. Gerthsen [BHS^{+}56]. Geschachtelte [Bau08c]. Geschichte [Bau90c, Bau00, Bau00f]. gesehen [BW72a]. gest [GB71, GB81, GB81]. gewölbter [BG51]. given [Bau76b, Ghi10]. Gleichungen [Bau54f]. Gleichungssysteme [Bau66a]. gleitendem [SB53]. Gnatz [Ghe87]. Goethe [BBG^{+}58]. Gomory [Bau63a]. gon [BH08]. Goos [Dij71, Dij72, Kad76]. Graeffe [Bau58a]. Graphs [Bau74b]. Gray [BBG^{+}58]. greater [Bau88a]. Green [Čul72b, Čul72a, Gep66]. Gregory [Bau07g]. Griechische [BBG^{+}58]. Gries [Ion84]. Grosse [MBH^{+}56]. Group [Ghe87, Bau52d, Bau52a]. Group-theoretical [Bau52a]. groups [Bau53, Bau54d, Bau54e, Bau55d, SBET85]. Grundlagen [Bau95a, BHS^{+}56, MBH^{+}56]. Grundlehren [Sch69, Sch70]. Gruppentheoretische [Bau52d, Bau52a]. guess [Bau03d]. Guggenheim [BBG^{+}58]. guided [BMPP89].

H [BW81, BH87, Bau95a, Čul72b, Čul72a, G.70, Gep66, Ghe87, Leh87, Sch69, Sch70, SBET85]. H. [BW81]. halbzahligem [Bau49a]. half [Bau49a]. half-integer [Bau49a]. Hamming [Bau07i, Bau07i]. Handbook [dLB07, BBG^{+}58, Sch69, Sch70]. Handbook/Guggenheim [BBG^{+}58]. Hangel [Ghe87]. Hartog [Bau09b]. Hartogs [Bau97e, Bau09b, Bau97e]. Hauptabschnitts [Bau61e]. Heidelberg [Kad76, Leh87]. Heidelberger [Dij71]. Heinhold [Knö66]. Heinz [Rut13, Bau71, Bau05f, Bau10h, Rut13]. heißt [Bau74c]. held [BS76, BBS93, Ghi10]. Herausgegeben [BBG^{+}63c]. Hesse [Ghe87]. Hg [Sch70]. hierarchies [BS76]. High [Rut13]. higher [Bau56a]. Hilbert [BBG^{+}58]. Hilfsmittel [AAB^{+}68, G.70]. Hill [LMNS16]. Hindenburg [Bau04e, Bau04e]. Hinsley [SBET85]. historian [Pet99].

Historic [BD01]. Historical [Bau74a, Bau77b, Bau09d]. Historiker [Pet99]. Historische [Bau09d]. History [Bau99b, MHR80, RH00, dLB07, AB87, Bau90c, Bau00, Bau00f, KHFA67]. Hochalpen [Rut13]. Hochschule [Bau11c]. höherer [Bau56a]. holds [Bau88a]. Hönger [BBG^{+}58]. honor [Ano05]. Hopf [Bau65b]. Hopfen [Bau74d]. Hopferau [Bau03a, Bau03a]. Horning [Ion84]. Householder [Sch69, Sch70, MB96]. Hurwitz [Bau55c, Bau55c]. Hurwitz-Zerlegung [Bau55c]. Hüttenhain [Bau08b, Bau08b]. Hydro [MBH^{+}56]. hypercubes [Bau91b]. Hyperwürfel [Bau91b].

ibid [Čul72a]. IFIP [FGR72]. Ihre [SBE^{+}99b, Bau08c]. II [Bau56c, Bau56b, Bau09f, Bau56c, BG71b, BGH76, BEH^{+}87, Bau09f, Kad76]. III [AAB^{+}68, G.70]. Im [Bau05c, Bau06d, Bau00d, Bau03a, SBE^{+}99b]. Imperial [Bau76b]. implementation [Bau79b]. Importance [Bau63b]. improved [Bau03d, HB62]. improvements [LMNS16]. Improving [Lev89]. impulses [Bau59b]. Inclusion [Bau66b]. Industriellen [BHS^{+}56]. inequalities [BH60b]. inequality [Bau61b, Bau65b]. infinite [Bau79a]. Influence [Bau61c, Bau61c]. Informatica [BB03]. Informatics [Bau88b, Bau09c, Bau10e, Bau07e, Bro15, Bau88b, Dij88]. Informatik [Bau74c, Bau88b, Bau90c, Bau00, Bau00f, Bau07e, Bau08d, SBE^{+}99b, BG71a, BG71b, BG73, BG74, BGH75, BGH76, BG91, BG92, Bau95a, Bau99c, Bau09d, Dij71, Dij72, Dij88, Kad76]. Informatik. [BG82, BG84]. Informatika [BG90, BG76]. Information [Bau00f, dLB07, BGH81, Hof62, FGR72]. Informationsverarbeitung [Hof62]. Informationswandler [Hof62]. informatyki [BGH81]. Ingenieurs [AAB^{+}68, G.70]. insbesondere [Bau56d]. Institut [Bau88b, Dij88]. Institute [BBS93, BBG^{+}58, Bau88b, Dij88]. integer [Bau49a]. integral [Bau03d]. Integralgleichungen [GTB53]. Integralgleichungen/Peters [GTB53]. Integration [Bau61a, Bau61d]. Intelligence [SBET85, CCP96]. interfaces [BS76]. interference [Bau58b]. International [BB79, BS76, BS00, CCP96]. Internazionale [Ghi10]. interview [AB87, SBE^{+}99b]. interviews [SBE^{+}99b]. Introduction [BFBS64, BHSS65, Rhe65, Car66, Hoh64]. Introductory [BG73, BG74, BG91, BG92, BG84]. intuition [BMPP89]. intuition-guided [BMPP89]. Intuitionism [Bau99c]. Intuitionismus [Bau99c]. invariance [Bau68b]. invented [Bau98b]. inventory [KHFA67]. inverse [Bau90b]. Inversion [BR71]. involving [BH60b]. Ionized [BBG^{+}58]. ISBN [Bau95a, Leh87]. isn’t [Knu90]. Issue [Ano05]. issues [Ghi10]. Italian [Ghi10]. Italy [Ghi10]. Iteration [Bau56d, Bau58a, Bau54f, Bau55c, Bau55a, BS57b, Bau57c, Bau57b, Rut69]. Iterationsverfahren [Bau54f, Bau55c, Bau57b, BS57b]. Iterationsverfahrens [Bau55a]. Iterative [HB60, RB55, RB55].

J [Bau64, Čul72b, Čul72a, G.70, Gep66, Ion84, Knö66, Sch69, Sch70]. J.

[BH87, Čul72b, Čul72a]. Jacobian [Bau54b]. Jacobischen [Bau54b]. Jahre [BW72a, Bau88b, Bau07e, Dij88, Bau85a]. jedermann [Hei68]. Jewish [Bau97e, Bau09b]. John [Bau07h]. Joint [Pry92]. Jordan [BR71]. Josef [Ale06]. Journal [Čul72a]. jüdischen [Bau09b]. jüdischer [Bau97e]. July [BS76, Bau91a, BBS93, BS00]. June [BD01].

Kalorimetrische [BBG^{+}58]. Kantorovic [Bau61b]. Kaplan [BHS^{+}56]. Karanapaddhati [Bau95b]. Karlson [Hei68]. Katz [Čul72b, Čul72a, Gep66]. kernels [BS57b]. Kerner [Gep66]. Kernsplitterung [BHB50]. Kettenbruch [Bau05b, Bau09e, Bau09f]. Kettenbruch-Phänomene [Bau09e, Bau09f]. Kettenbrüche [Bau06e, Bau09e]. Kettenmatrix [Bau55b]. Kettenprodukte [Bau08e]. key [HB62]. Kfoury [Ion84]. Kinematographie [BBG^{+}58]. Klaus [BEPP81, GB81, GB81]. kodierten [BS57a, BS60b, Bau01b]. Koecher [Bau89a]. Koechers [Bau89a]. Köhler [BHS^{+}56]. Komma [Bau96d, SB53]. Kommission [Bau00, Bau00f]. Kompendium [BBG^{+}58]. Kompressible [BG51]. Konrad [Roj96, BW72b, Bau85a, Bau96b, Bau00e, Bau03a, Bru11, Roj96]. Konstante [Bau89a]. konvergente [Bau54b, Bau54f, Bau55a]. Konvergenz [Bau54a]. Krige [Ghe87]. Krige-Brückner [Ghe87]. Kristalle [BHS^{+}56]. Kristalle/Przibram [BHS^{+}56]. Kristallphotozellen [MBH^{+}56]. Kristallphotozellen/Czech [MBH^{+}56]. Kryptologie [Bau00c, Bau82b, Bau93, Bau94b, Bau95c, Bau96c, Bau97c]. Kudrjawzew [BHS^{+}56]. Kunzi [G.70]. kurs [BG76]. kurzer [SB54, Bau90c]. Kurzes [BBG^{+}58]. Kybernetik [Bau65a, LB69].

L [Ale06, Ano04, Ano05, Ano16, Ano19, AB87, BGHS79, Bod04, Bro15, Car66, Chr07, Chu51, Čul72b, Čul72a, DS76, Dij86, G.70, Gan15, Gep66, Ghe87, Ghi74, Hei68, Hoh64, Ion84, Kad76, Kil12, Knö66, Kob10, Leh87, Liv81, Pet99, Rhe65, Rut69, Sch69, Sch70, Scr79, SBE^{+}99a, Ghe87]. L.N.C.S [Ghe87]. Lagrange [Bau05a, Bau05a]. Laguerre [Bau03d]. Lambert [Bau05b]. Lamberts [Bau05b]. Language [BBG^{+}63c, BBG^{+}76, BBR^{+}58, BBG^{+}78a, BBD^{+}81, BW82, BB84, Ghe87, NBB^{+}60, NBB^{+}63, NBB^{+}97, vWMPK69b, vWMPK69a, BBG^{+}60, BBG^{+}63a, BBG^{+}63b, BS60a, Bau78, BBG^{+}78b, BBG^{+}79, BW81, Čul72b, Čul72a, EPBS63, Gep66, NBB^{+}62, NBB^{+}65, BS76]. Languages [Bau10c, BS62b, Bau66c, BW72b, BW72a, Bau97d, HB62]. Laut [Ghe87]. law [OML09]. lazy [Bau79a]. Least [Bau65c]. Lebensgeschichte [SBE^{+}99b]. Lebenswerk [Leh87]. Lect [Liv81]. lectures [Bau76b, Bau68c, Ghi10]. Legenden [Bau96b]. Legendre [Bau03d]. legends [Bau96b]. Lehrbuch [BBG^{+}58]. Leibniz [Bau07g, Bau07g]. Leinen [Hei68]. Leitfäden [Knö66]. Leonhard [Bau07c, Bau07g, Bau07c, Bau07g]. Letter [Bau95b]. Letters [AGG60, LYB98, RBHM95]. Liebmann [Bau97e]. Life [Gan15, Bro15, SBE^{+}99b]. Limits [Bau97a]. Linear [Bau65c, HB60, WR71, Bau55a, Bau57b, Bau66a, BDS69]. linearen [Bau55a, Bau57b]. linearer [Bau66a, BDS69]. List [HB62]. Ljapunow [Bau65a]. Ljubjana [FGR72]. logarithmic [Bau03d]. Logic [Bau87, Bau91a, BBS93, Bau49b, BW91, LB69]. Logical [Bau60a]. Logik [LB69]. Logikkalküls [Bau49b, Chu51]. London [Bau76b]. Lösung [Bau54f, Bau57c, Bau66a]. Lösungen [BGH75, BGH76]. Louis [Bau54c]. Ludwig [FGH^{+}06]. Lumineszenz [BHS^{+}56]. Lumineszenz/Smyth [BHS^{+}56].

M [Car66, Čul72b, Čul72a, Gep66, Ghe87, Hoh64, Ion84, Liv81, Rhe65, SBET85, Bau65a]. M. [Bau54c]. machine [BS57a, BS60b, BS62b]. Machinery [Čul72b]. Machines [BS62a, Bau96a, Bau10c, Bau01b, Bau07b, Roj96]. macro [OML09]. made [Bau03d]. Madison [Lan59]. Maehly [BS62c]. magic [Bau04b, Bau96e, Bau04b]. magische [Bau04b, Bau04b]. Magnetic [Bau50]. Magnetische [Bau50]. main [Bau61e]. mapping [Šve70]. Marian [Bau05c, Bau06d, Bau00d, Bau05c, Bau06d]. Marktoberdorf [BS76, Bau91a, BBS93, BS00, BB03]. Marquartstein [KB53]. Marsprojekt [GTB53]. Marsprojekt/Neher [GTB53]. Maschinelle [BS62b]. Maßnahmen [SB54]. Matematico [Ghi10]. MATHEMATICA [BH08]. Mathematical [BBG^{+}58, GTB53, AAB^{+}68, CCP96]. mathematician [Bau09b]. mathematicians [Bau97e]. mathematics [Bau68c, Bau92, Bau05c, Bau06d, EK81, SB90, Sto93, Sto94]. Mathematik [Ale06, Bau92, Čul72b, Čul72a, Hei68, Knö66, Bau68c, Bau92, Bau05c, Bau06d, SB90, Sto93, Sto94]. Mathematiker [Bau97e]. Mathematikers [Bau09b]. Mathematische [AAB^{+}68, G.70]. mathematischen [Sch69, Sch70]. matrice [RB55]. Matrices [Bau63c, Bau63b, Bau69, BR71, Pry92, RB68, Bau52c, Bau54c, Bau65e, Šve70, Wil84]. Matrix [Bau58a, wC03, Bau54c, Bau55b, BH60b, Bau61e, wCL04, NS64, RB55, Zen68, Bau61e]. Matrizen [Bau65e]. maturity [Dij88]. Matzner [Ghe87]. Maxima [Bau00c, Bau95c]. Maximen [Bau95c, Bau97c, Bau00c, Bau82b, Bau93, Bau94b]. Maxims [Bau97b, Bau00a, Bau00b, Bau02a, Bau07d, Chr07, Kob10, Bau82b, Bau93, Bau94b, Bau97c]. Mayer [BHS^{+}56]. McCarthy [Čul72b, Čul72a, Gep66]. MDN [Bau65a, Gep66]. mean [Bau74c]. Measures [SB54]. Mechanical [Bau10b, BB49]. mechanics [Che72]. Mechanik [Knö66]. Mechanisierung [KB53]. mechanism [Bau77a]. Mechanization [KB53]. Meetings [MB96]. Mein [Leh87]. Memorial [BS81]. memoriam [BH87, Bau89b, Bau07h]. Memories [BS81]. memory [BBG^{+}63c, Bau05c, Bau06d]. Menschen [GTB53]. messages [Bau53, Bau54d, Bau55d, Bau58b]. Messtechnik [BBG^{+}58]. Messtechnik/Hönger [BBG^{+}58]. Methke [MBH^{+}56]. Method [Bau58b, Bau59b, BS62a, BR71, Bau54b, Bau54a, Bau54c, Bau55c, Bau56d, Bau57c, Bau61d, Lev89, RB55, Rut69, BS57a, BS60b, Bau01b]. Methode [Bau54b, Bau54c, Bau61d, RB55, Bau54a]. Methoden [Bau93, Bau94b, Bau95c, Bau97c, Bau00c, BBG^{+}58]. Methodos [Chu51]. Methods [Bau97b, Bau00a, Bau00b, Bau00c, Bau02a, Bau07d, Chr07, HB59, HB60, Kob10, Bau54b, Bau54f, Bau56c, BS57b, Bau57c, Bau57b, Bau93, Bau94b, Bau95c, Bau97c, BBG^{+}58, MBH^{+}56]. Michael [Ion84]. micro [OML09]. micro-macro [OML09]. Microelectronics [Bau10e]. Might [Bau97a]. Millward [LMNS16]. Minimal [Bau63b]. Mitarb [Gep66]. Mitarbeit [BW81]. Mittel [BBG^{+}58]. model [Bau51]. Modern [Bau58a, EK81, BHSS65]. Moderne [Knö66, BHSS65]. Modified [BBG^{+}76]. Moivre [Bau05a, Bau05a]. Moll [Ion84]. Möller [Ghe87]. Moments [BH60a]. Momentum [MBH^{+}56]. Momentum/Zimmer [MBH^{+}56]. Monographs [Ion84]. Monotonic [BSW61, Bau90b]. Morse [MBH^{+}56]. most [BHB50]. motives [BS60a]. Mr [SBET85]. Muir [BKB13]. Müller [BG51]. multiple [Bau05a]. Multiplication [Bau94a]. München [Bau07e, Bau97e, Bau09b, Bau11c]. Munich [Bau74d, Ghe87, BBB^{+}85, BEH^{+}87, Bau96a, Bau97e, Bau07e, Bau09b, Bau11c]. murmur [Bau96e]. Museum [Bau96a]. MuseumsForum [Bau10h]. Mutter [BBG^{+}58]. Mutually [Bau10e]. My [Bau02b, Bau03b, Bau11b].

N [Ion84]. NATO [BBS93]. natural [Bau66c]. natürlicher [Bau66c]. Naturwissenschaft [BBG^{+}58]. Naturwissenschaft/Erckmann [BBG^{+}58]. Naur [Čul72b, Čul72a, Gep66, AGG60, BBG^{+}63c]. near [Bau74d]. Neher [GTB53]. neighbouring [Wil84]. Nested [Bau08c]. Neu [Hei68]. Neumann [Bau98b, Bau98b]. Neutronen [BHB50]. neutrons [BHB50]. Newton [Bau55a, Bau55a, BS62c, RB68]. Newton-Prozess [Bau55a]. nichtsymmetrischer [Bau65e]. Nickl [Ghe87]. Niederfrequenz [BBG^{+}58]. Nixdorf [Bau10h]. No [BBG^{+}63c, Čul72a]. non [Bau65e, Bau90b]. non-monotonic [Bau90b]. non-symmetrical [Bau65e]. nondominant [Lev89]. Nonlinear [Bau65d]. norm [Bau62]. normal [Bau61e]. normalen [Bau61e]. Normen [BD64]. Norms [BF60, BSW61, Bau75, BH61]. Note [BF58]. Notes [Liv81, Bau09d]. Notizen [Bau09d]. Nova [Bau05d]. nuclear [BHB50, BHS^{+}56]. nucleus [Bau51]. Nullstellenbestimmung [Bau54b, Bau56d]. number [Šve70]. Numbers [Bau09f]. numerica [Ghi10]. Numerical [Bau74d, Bau10f, Bau54b, Bau56c, Bau56a, BS59, BS60a, Bau61d, BRS63, Bau68c, Ghi10, Lan59, Bau65e, SB90, Sto93, Sto94]. Numerics [Rut13]. Numerik [Rut13]. numérique [Bau61d]. Numerische [Čul72b, Čul72a, Bau68c, Ale06, Bau65e, SB90, Sto93, Sto94]. numerischen [Bau56a, BS59]. numerischer [Bau54b, Bau56c].

O [Gep66]. O. [SBET85, DS76]. obituary [Bro15]. objects [Bau79a]. occasion [BS81, Bau07c, Bau07g]. October [Bau76b]. odd [Bau08a]. Oedometric [OML09]. Oh [Wei88]. old [Bau85a]. Olga [Bau98a]. Oliver [Sch70]. Olver [Sch69]. One [Gri90, Bau98c]. one-step [Bau98c]. Operation [BS62a]. operational [Bau55b]. Operatoren [BDS69]. operators [Bau65b, BDS69]. Optimal [Bau63b, SB53]. Optimale [SB53].

Optimally [Bau63c, Bau69]. Oral [AB87]. order [Bau55a, Bau56a, Lev89]. Ordnung [Bau55a, Bau56a]. Original [BD01]. Origins [Bau10h, Kil12]. Oszillograf/Reichardt [MBH^{+}56]. our [FvGGM90]. outline [Bau90c]. Overview [BG91, BG92, BG84].

P [BBG^{+}63c, BW81, Čul72b, Čul72a, G.70, Gep66, Ghe87, Hei68]. P. [BBG^{+}63c, Bau64, BW81]. Packings [SBET85]. pairs [wCL04]. Pake [MBH^{+}56]. Papers [Ano05]. parenthesis [Bau77a]. parenthesis-free [Bau77a]. Part [Sch70, BG73, BG74, BGH75, BG82, BG84, AAB^{+}68, BGH76, Bau05c, Bau06d, Bau09e, Bau09f]. particles [Bau49a, BHB50, BB49]. particles/nuclear [BHB50]. particular [Bau56d]. Partsch [BW81, Ghe87, BW81]. Pepper [Ghe87, BW81]. Period [Bau96d]. Periodic [Bau06e, Bau09e, wCL04]. Periodische [Bau06e, Bau09e]. Perlis [Čul72b, Čul72a, Gep66]. Perturbation [wC03, wCL04]. Perugia [Ghi10]. Peter [AGG60]. Peters [GTB53]. Petrovich [Bau89b]. Petrovskij [GTB53]. Phänomene [Bau09e, Bau09f]. Phenomena [Bau09e, Bau09f]. philosophy [Bau76a, Bau76b]. Photoelemente [MBH^{+}56]. Photographie [BBG^{+}58]. Photographie/Photographie [BBG^{+}58]. physical [Che74, KB53]. Physics [BBG^{+}58, KHFA67, BHS^{+}56, BBG^{+}58, MBH^{+}56]. Physics/Bechert [BHS^{+}56]. Physics/Blaschke [BBG^{+}58]. Physics/Feenberg [MBH^{+}56]. Physics/Mayer [BHS^{+}56]. Physik [BG51, BHS^{+}56]. Physik/Frössel [BG51]. physikalischchemischen [BHS^{+}56]. physikalische [KB53, BBG^{+}58]. Physikalischen [BBG^{+}58]. Pioneer [Gan15]. Pioneers [BD01, Wei88, BD01, SBE^{+}99b]. Pioniere [SBE^{+}99b]. pivot [Bau61c]. Planck [BBG^{+}58]. Planeten [GTB53]. Plankalkül [BW72b, BW72a, BW72a, Bau00e, Bau03a]. Poincaré [Lev89]. Poincaré-type [Lev89]. Point [Bau50, Bau99a, SB53]. pointer [Bau79a]. Poland [Bau05c, Bau06d]. Polen [Bau05c, Bau06d]. Poles [Bau50]. Polish [BGH81]. polygons [Bau91b]. Polygraphia [Bau05d]. Polynomen [Bau54b]. Polynomial [Bau56b, HB59, Bau55c, Bau56c, Bau56d, BS57b]. Polynomials [wC03, Bau54b, Bau66b]. Polynomkerne [BS57b]. Polynoms [Bau55c, Bau56c, Bau56b, Bau56d]. Positive [BR71, Bau65b, BDS69]. positiver [BDS69]. Positivity [Bau75]. pp [Chu51, Čul72b, Čul72a, G.70]. Preamble [Bau10g]. Preis [Bau65a, Gep66]. Preliminary [AGG60, Bau05c, Bau06d]. Price [G.70]. primes [Bau88a]. Principle [Bau90a]. Pringsheim [Bau97e]. problem [BS60a, Bau90b]. Probleme [Bau57b, Hof62, Bau65a]. Problèmes [BBG^{+}58]. Problèmes/Roth [BBG^{+}58]. Problems [Bau65c, Bau54f, Bau56a, Bau56d, Bau57c, Bau57b, BGH81, Hof62]. procedures [Bau82b]. proceedings [EK81, FGR72, Bau91a, BBS93, Lan59, CCP96]. Process [Bau53, Bau54d, Bau55d, Bau55a, Bau57a, Bau76c, Bau76d]. Processes [Bau58a, BBR^{+}58]. processing [BS57a, BS60b, BS62b, Bau01b, FGR72, Hof62]. processors [EPBS63, Hof62]. products [Bau08e]. Prof [DS76]. Prof. [Bro15]. Program [BBG^{+}78a, BB79, Bau79c, BW82, Bau54b, Bau56c, Bau76e, Bau78, BBG^{+}78b, BBG^{+}79, BW81, BEH^{+}87, BMPP89, Knu90, Liv81]. program-controlled [Bau54b, Bau56c]. Programme [SB54]. Programmentwicklung [BW81]. programmgesteuerte [Bau54b, Bau56c]. Programmiersprachen [BW72a, BS62b]. Programmierung [HB62]. Programmierungstechnik [HB62]. Programming [Bau76c, Bau76d, BBD^{+}81, Ion84, BS62b, BW72b, BW72a, Bau76a, Bau76b, Bau78, BMPP89, HB62, PPWB78]. Programmtransformation [Bau76e]. Programs [BB84, SB54]. Project [BBB^{+}85, BEH^{+}87, Ghe87, Bau79c, SB62]. Promotionsrecht [Bau11c]. proof [Bau65b, Knu90]. properties [BF58, Zen68]. Proposal [BBR^{+}58]. propositional [BW91]. propres [RB55]. Protonen [BHB50]. protons [BHB50]. Prozess [Bau55a]. Przibram [BHS^{+}56]. pulse [Bau53, Bau54d, Bau55d]. Punkt [Bau96d, Bau99a]. Punktpole [Bau50]. Pythagoräische [Bau05e]. Pythagorean [Bau05e].

Quadrate [Bau04b]. quadratic [Bau55a, RB55, Wil84, Bau54a]. quadratically [Bau54f, Bau54b]. quadratique [RB55]. Quadratisch [Bau54b, Bau54f, Bau55a]. Quadratische [Bau54a]. quadrature [BRS63]. quadrupoles [Bau55b]. Quantenphysik [MBH^{+}56]. Quantentheorie [BHB50]. quantum [BHB50, Che72, KHFA67, BHS^{+}56]. questioni [Ghi10]. Questions [Bau66a]. quotient [Bau59c]. QWERTY [Bau03c]. QWERTZU [Bau03c].

R [Car66, G.70, Ghe87, Hoh64, Ion84, Knö66]. R. [G.70, Rhe65]. Raaz [BHS^{+}56]. rang [Bau54c]. rank [Bau54c]. Raster [Bau07f]. Rational [RB68, Bau05a]. rationalen [Bau05a]. Reaktionswärmen [BBG^{+}58]. Reaktionswärmen/Kurzes [BBG^{+}58]. Rechenanlagen [Bau54b, Bau56c, BHSS65, HB62, Knö66, SB53]. Rechenautomaten [SB54]. Rechengenauigkeit [SB53]. Rechenmaschine [BS57a, BS60b]. Rechenmaschinen [Bau01b, Roj96]. Rechenstrukturen [PBB^{+}82]. Rechentechnik [Bau57b]. Rechner [Bau98b]. recognition [Bau92]. rectorate [Bau11c]. Recueil [BBG^{+}58]. recurrence [Lev89]. Reddy [Ion84]. Reduced [Bau49a]. Reduction [Bau59a]. Regia [BBG^{+}58]. Regia/Courant [BBG^{+}58]. regular [Bau06a]. Reichardt [MBH^{+}56]. Reihen [Bau89a]. Rejewski [Bau00d, Bau05c, Bau06d, Bau00d, Bau05c, Bau06d]. Rektoratsverfassung [Bau11c]. Related [Bau65f, Bau10a, Bau57c]. Relationen [Bau07g]. relations [Bau07g, Lev89]. Relationships [Bau57b]. Remark [Bau07a, Bau89a, Bau06a]. Remarks [Bau69, Bau74a, Bau77b]. Remembers [MB96]. Removal [Bau09f]. removing [Bau58b]. Reply [Bau64]. Report [AGG60, BBG^{+}60, BBG^{+}76, NBB^{+}60, NBB^{+}63, NBB^{+}97, vWMPK69b, vWMPK69a, BBG^{+}63a, BBG^{+}63b, Čul72a, KHFA67, NBB^{+}65, BBG^{+}63c, Čul72b, Gep66, NBB^{+}62]. representation [Bau79a]. representations [Bau52b, Bau52b]. Reprint [SB83]. restrictions [BS60a]. results [BS60a]. Reuber [BBG^{+}58]. Review [Ale06, BG73, BG74, Car66, Chr07, Chu51, Čul72b, Čul72a, Dij71, G.70, Ghe87, Ghi74, Hoh64, Ion84, Kil12, Knö66, Kob10, Leh87, Liv81, Rhe65, Sch69, Sch70, Scr79, Dij72]. Reviews [Bau95a, SBET85]. Revised [BBG^{+}63a, BBG^{+}63b, NBB^{+}63, NBB^{+}65, NBB^{+}97, Čul72a]. revisited [Bau00e]. Richard [Bau07i]. Right [Bau11c]. Robert [BS81, Ion84, BS81, GB71]. Rocket [MBH^{+}56]. Roland [Ale06]. Romberg [Bau61a, Bau61d]. Roots [BH60a, Bau54b, Bau56d, BH61, Bau66b, Bau08c]. Roth [BBG^{+}58]. Rotor [Bau99b, Bau07b]. Rounding [Bau74b, Bau61c]. Row [Bau65c]. rules [BBP^{+}79]. Rundfunkröhren [BBG^{+}58]. Rundfunkröhren/Mutter [BBG^{+}58]. Russian [Bau82a, NBB^{+}65]. Rutishauser [Čul72b, Čul72a, G.70, Gep66, Sch69, Sch70, Bau71, Bau80, Bau02b, Bau03b, Bau05f, Bau11b, Rut13].

S [Bau65a, Bau95a, Gep66, Hei68, Ion84, Kad76, Leh87, Sch69, Sch70, BEH^{+}87]. Sackgassen [Bau04c]. Salute [FvGGM90]. Samelson [Car66, Čul72b, Čul72a, G.70, Gep66, Ghe87, GB81, GB81, Knö66, Rhe65, Sch69, Sch70, BEPP81, GB81, Hoh64]. Samet [Bau64]. sand [OML09]. Sauer [BS81, G.70, GB71, Knö66, BS81]. Scaled [Bau63c, Bau69, Šve70]. Scaling [Bau63b, Bau68b]. scattering [BHB50]. Schalenmodell [Bau51]. Scheinlichkeitsrechnung [BBG^{+}58]. Scheinlichkeitsrechnung/Gray [BBG^{+}58]. Schichten [BHS^{+}56]. Schichten/Raaz [BHS^{+}56]. Schicksal [Bau09b]. Schicksale [Bau97e]. School [BB79, Ghi10, BS76, BS00]. Schriften [BBG^{+}58]. Schulz-Methke [MBH^{+}56]. Schweizer [Rut13]. Science [BG73, BG74, Bau76b, BG91, BG92, Bau00, Bau00f, Bau08d, Ion84, Liv81, Bau74c, BGH75, BGH76, BGH81, BG82, BG84, Bau99c, Bau09d, Dij88, EK81, SBE^{+}99b]. Sciences [Bau00, Bau00f]. scientific [BS60a]. sd&m [BD01]. se [Bau54c]. Sears [BHS^{+}56]. Seaton [MBH^{+}56]. Sechsstellige [GTB53]. Second [SBET85, Lev89, BG74, BG84, HB62]. second-order [Lev89]. Secret [SBET85]. Secrets [Bau97b, Bau00b, Bau00c, Bau02a, Bau07d, Bau09a, Chr07, Kob10, Bau95c, Bau97c, Bau00a]. section [Bau61e]. secure [BS00]. securing [Bau53, Bau54d, Bau55d, Bau59b]. Security [dLB07]. See [Bau74d]. seen [BW72a]. seines [BS81]. Seit [Bau06e]. selected [Hof62]. September [CCP96, EK81, Ghi10]. sequence [Bau65d]. Sequential [Bau59a, SB60, SB83, SB59, SB01]. Sequentielle [SB59, SB01]. series [Bau89a, Bau95b]. sets [Bau66b]. Shannon [Bau01a, Bau01a]. Shape [Bau09c]. sharp [SW12]. shell [Bau51]. Shift [RB68]. short [SB54]. Simon [Bau06c]. Simon‘ [Bau06c]. Simple [Bau06c, SBET85, Knu90]. simplexes [Che72, Che74]. simultaneous [Rut69]. Since [Bau06e]. Smyth [BHS^{+}56]. Software [Bau72, Bau73a, Bau73b, BDW^{+}75, Bau77c, BD01, Bau06b, Ghi74, Scr79]. software“ [Bau06b]. solution [Bau90b]. Solutions [Bau10c, BGH75, BGH76, Lev89]. Solving [Bau65c, HB60, Bau54f, Bau57c, Bau66a]. Some [BH60b, Bau68b, Bau57b, Ghi10]. Sommerfeld [BG51, GTB53]. sont [Bau52c]. Sources [KHFA67]. Soviet [Bau07j]. Sowjetunion [Bau07j]. Special [Ano05, WPPB79]. Specification [BBG^{+}78a, BBG^{+}78b, Bau79b, BBG^{+}79, BBS93, PBB^{+}82]. specifications [BW88]. Spectrum [BBG^{+}78a, BBD^{+}81, Ghe87, Pry92, BBG^{+}78b, BBG^{+}79]. Speiser [Bau08d]. Spezifikation [PBB^{+}82]. Sphere [SBET85]. spin [Bau49a, Bau52d, Bau52a, Bau54e, Bau49a]. Spingruppen [Bau54e]. spinor [Bau52b]. spinorielles [Bau52b]. spins [BB49, BB49]. Spinwellengleichungen [Bau52d, Bau52a]. Spitzer [BBG^{+}58]. Sprache [BBG^{+}63c, BW81, Gep66, NBB^{+}62]. Sprachen [HB62, Bau66c]. Springer [Bau95a, Dij71, G.70, Kad76, Leh87]. Springer-Verlag [Bau95a, G.70, Leh87]. Squares [Bau65c, Bau04b]. SSR [EK81]. Stabler [GTB53]. standards [BD64]. Stanislaus [Bau60a]. Star [Bau91b]. State [Bau90a]. states [Che72, Che74]. stationären [Bau55a]. stationary [Bau55a]. statistical [Che72]. Statistik [GTB53]. Statistik/Braun [GTB53]. step [Bau57c, Bau98c]. stepwise [Bau79c]. Sternpolygone [Bau91b]. Steyr [CCP96]. Stichworten [HB62]. Stiefel [BGHS79, Sch69, Sch70]. Stirling [Bau07a]. Stock [Bau61e]. Stoer [Ale06, G.70]. Storage [Bau90a]. stories [SBE^{+}99b]. Streuversuche [BHB50]. Strömung [BG51]. Structure [Che72, Bau66c, BHS^{+}56, Bau92]. Structure/Kudrjawzew [BHS^{+}56]. structures [PBB^{+}82]. Struktur [Bau66c]. Strukturerkennung [Bau92]. studies [Bau52d, Bau52a]. Study [BBS93]. style [Bau08e]. subordinate [Bau62]. Summer [BB79, Ghi10, BS76, BS00]. Supplement [BBG^{+}63c]. Support [BBG^{+}78a, BBG^{+}78b, BBG^{+}79]. Swiss [Rut13]. symbolic [CCP96]. Symbols [Bau10d]. Symmetric [RB68]. symmetrical [Bau65e]. Symposium [Bau74d, Lan59]. Syntactic [Bau66c]. Syntaktische [Bau66c]. syntax [EPBS63]. system [Bau78, BEH^{+}87, Bau94a]. Systematic [BB84]. Systematics [BBP^{+}79]. Systems [HB60, Bau54b, Bau56c, BHSS65, Bau66a, Che74]. Szabo [G.70].

T [G.70, Ghe87]. T. [SBET85]. Tafel [GTB53]. Take [Bau09c]. Taschenbücher [Dij71]. Taussky [Bau98a]. Taussky-Todd [Bau98a]. Taylor [Bau64]. Technical [Bau11c, HB62]. techniques [WPPB79]. Technischen [Bau11c]. Technology [Bau76b, Bau57b, HB62]. Teil [AAB^{+}68, BG71a, BG71b, BGH76, Bau05c, Bau06d, Bau09e, Bau09f, BG73, BG74, BGH75, BG82, BG84, Dij71, Dij72]. telegraph [Bau59b]. telegraphic [Bau58b]. Tenseurs [Bau52c]. Tensors [Bau52c]. terms [HB62]. ternärer [Bau07j]. Ternary [Bau07j]. test [OML09]. testimony [Bau06b]. Texts [Ion84]. Their [Bau97a, BD01, Bau08c, SBE^{+}99b]. Them [Bau65f, Bau10a].

Theodor [Bau04d]. Theorem [Pry92, Li88, SW12, wEC86, Dij86]. Theorems [BF60, wC03, wCL04]. Theoretical [MBH^{+}56, Bau52d, Bau52a]. Theoretical-physical [KB53]. Theoretisch-physikalische [KB53]. Theoretische [BG51]. Theorie [Bau52d, Bau52a, Bau54e]. Theory [BHS^{+}56, BHB50, Bau52d, Bau52a, Bau54e]. Thermodynamics [BBG^{+}58]. Thermodynamics/Bruhat [BBG^{+}58]. Thermodynamik [GTB53]. Thilo [BBG^{+}58]. Thompson [SBET85]. Thought [GTB53]. Thought/Petrovskij [GTB53]. Thrall [Bau10b]. Time [Gri90]. Tl [Kad76]. today [BW72b]. Todd [Bau98a, Bau07h]. Todestages [BS81]. Tokyo [Leh87]. tools [AAB^{+}68]. Transcendental [Bau09f]. Transform [BKB13]. Transformation [RB68, BS59, Bau76e, Bau78, BBP^{+}79, BEH^{+}87, WPPB79, BS59]. transformational [PPWB78]. transformations [Bau65d, Bau79c, BMPP89]. transformations-computer-aided [BMPP89]. Transition [Bau90a]. Translation [GHL67, SB60, SB83, SB59, SB01]. transmission [Bau53, Bau54d, Bau55d, Bau59b]. transzendente [Bau09f]. treatment [Bau56a]. tree [Bau08a]. Treppeniteration [Bau57c]. Tridiagonal [Bau59a, RB68]. trigonometrischen [GTB53]. Tripel [Bau05e]. Triple [Bau05e]. Trits [Bau07j]. trouvant [Bau54c]. Trytes [Bau07j]. Tsichritzis [Ion84]. Turanski [BBG^{+}63c, BBG^{+}63c]. Twentieth [MHR80]. Twenty [Bau88b, Dij88]. Type [Bau58a, Li88, Lev89]. types [Bau85b, PBB^{+}82]. typographical [Bau99a]. typographische [Bau99a].

u [Gep66]. Übers [Gep66]. Übersicht [BG71a, BG71b, BG73, BG74, BG84, Kad76, BG91, BG92]. übersichtlicher [SB54]. Übertrag [Bau98c]. Ullstein [Hei68]. Ultraschallverfahren [BHS^{+}56]. ungerade [Bau08a]. Union [Bau07j]. Universal [BBR^{+}58, Bau10c]. Universalis [Bau05d]. University [Bau76b, Bau11c, BHS^{+}56]. unreasonably [Bau05c, Bau06d]. unterhaltsame [Hei68]. Untersuchungen [Bau52d, Bau52a, BHS^{+}56]. Untersuchungen/Davis [BHS^{+}56]. unvernünftig [Bau05c, Bau06d]. Upper [MBH^{+}56]. Urgench [EK81]. Use [Bau65f, Bau10a]. used [Bau05c, Bau06d]. Using [Bau10d]. Uzbek [EK81].

V [Ion84]. values [Bau61e, Bau62, Bau68a, NS64, Zen68]. Vauquois [Čul72b, Čul72a, Gep66]. vecteurs [RB55]. Verarbeitung [BS57a, BS60b, BS62b, Bau01b]. verbess [HB62]. Verfahren [Bau54b, Bau56c, Bau56d, Bau57c, Bau82b, Bau57a, BS57a, BS60b, Bau01b]. Verfahrens [BS57a, BS60b, Bau01b]. Verfärbung [BHS^{+}56]. Verlag [Bau65a, Bau95a, BBG^{+}58, Dij71, G.70, Gep66, Hei68, Kad76, Leh87]. verschiedenen [BB49]. Verschwörung [Bau11a, Bau12]. version [SW12]. verwandte [Bau57c]. VI [Bau74d, Gep66]. via [wC03, wCL04, BBG^{+}58]. Vielfachen [Bau05a]. Vierpolen [Bau55b]. vitae [FGH^{+}06]. Vol [Sch69, Sch70, Chu51, Čul72b, Čul72a, Ghe87]. Vol. [BBB^{+}85]. Volume [Ghe87, BEH^{+}87]. vom [BW72a]. von-Neumann-Rechner [Bau98b]. Vorarbeiten [Bau05c, Bau06d]. Vorläufer [BW72a]. Vorlesungen [Ale06]. Vorträge [Bau68c, BBG^{+}58]. Vorträge/Goethe [BBG^{+}58]. vvodnyi [BG76].

W [BB03, Čul72b, Čul72a, FvGGM90, Gep66, Ghe87, Ion84, Sch70]. Wahr [BBG^{+}58]. Waite [Ion84]. Wallis [Bau08e]. Wallis-artige [Bau08e]. Wallis-style [Bau08e]. Wand [BG51]. Wand/Müller [BG51]. War [Bau04e, SBET85]. Warum [Bau85b]. was [Bau74c, Bau74c, Bau76e, Bau04e]. wave [Bau49a, Bau52d, Bau52a, BB49]. Wechselwegnahme [Bau09f]. Weed [BHS^{+}56]. Wegstein [Čul72b, Čul72a, Gep66]. Weighted [Bau65c]. Wellengleichungen [Bau49a, BB49]. wellformedness [Bau77a]. Wer [Bau98b]. Wertevorrat [Bau61e]. Who [Bau98b]. whose [Bau52c, Knu90]. Wide [BBG^{+}78a, BBD^{+}81, Ghe87, BBG^{+}78b, BBG^{+}79]. Wien [Bau95a]. Wien/New [Bau95a]. Wijngaarden [Čul72b, Čul72a]. Wijngarden [Gep66]. Wilkinson [BH87]. William [BBG^{+}63c, BBG^{+}63c]. Wirk [BBG^{+}58]. Wirsing [Ghe87]. Wissenschaften [Bau00, Bau00f, Sch69, Sch70]. Wittmeyer [Bau55a, Bau55a]. Wittmeyer-Prozess [Bau55a]. Woodger [Čul72b, Čul72a, Gep66]. words [HB62]. work [BS60a, Bau05c, Bau06d, Rut13]. working [KB53, KB53]. World [SBET85]. Wörterbuch [HB62]. Wörterverzeichnis [HB62]. Wössner [Ghe87]. wrong [Bau03d]. Würfel [Bau04b]. Wurzeln [Bau08c].

xiii [Bau95a]. xv [Leh87]. xviv [G.70].

year [BW72a, Bau03b]. years [Bau85a, Bau88b, Bau02b, Bau07e, Bau11b, Dij88]. York [Bau95a, Kad76, Leh87]. York-Tokyo [Leh87]. Yugoslavia [FGR72].

z [BGH81]. Z1 [Roj96, Roj96]. Z3 [Roj96, Roj96]. Z4 [Bau03a, Bau03a]. zadań [BGH81]. Zahlen [Hei68, Bau09f]. zahlr [Hei68]. Zauber [Hei68]. Zaubergemurmel [Bau96e]. Zbiór [BGH81]. Zemanek [Leh87]. Zemansky [BHS^{+}56]. Zerlegung [Bau55c]. Zeugnisse [Bau06b]. Zimmer [MBH^{+}56]. zum [Bau57a, BGH75, BS81]. Zur [BBG^{+}63c, BBG^{+}58, Rut13, Bau52d, Bau52a, Bau54b, Bau54f, Bau55c, Bau56c, Bau56d, BS57a, Bau57c, BS60b, Bau01b, Bau09d, SB54, Bau49b, BHB50, Bau54e, Bau56a, BS59, Bau00, Bau00f, Chu51]. Zurich [BS60a, Bru11, Bru11]. Zusammenhänge [Bau57b]. Zuse [Bau85a, Bau96b, Bau03a, Bru11, BW72b, Bau80, Bau85a, Bau96b, Bau98c, Bau00e, Bau03a, Bru11, Leh87, Roj96]. Zuses [BW72a]. Zwanzig [Bau88b, Dij88]. Zweiter [BG74, BG84, Dij72]. zwischen [Bau57b, GTB53].

References

Angelitch:1968:MHI

[AAB^{+}68] T. P. Angelitch, G. Aumann, F. L. Bauer, R. Bulirsch, H. P. Künzi, H. Rutishauser, K. Samelson, R. Sauer, and J. Stoer. Mathematische Hilfsmittel des Ingenieurs. Teil III. (German) [Engineer’s mathematical tools. Part III]. Herausgegeben von R. Sauer und I. Szabó. Unter Mitwirkung von H. Neuber, W. Nürnberg, K. Pöschl, E. Truckenbrodt und W. Zander. Die Grundlehren der mathematischen Wissenschaften, Band 141. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. ISBN 3-642-95030-2, 3-642-95031-0. xix + 535 pp. LCCN QA76 .B3192.

Aspray:1987:OHI

[AB87] William Aspray and Friedrich L. Bauer. Oral history interview with Friedrich L. Bauer. Web site at the Charles Babbage Institute, The Center for the History of Information Processing, University of Minnesota, Minneapolis, MN, USA., Febru-

ary 17, 1987. URL http://hdl.handle.net/11299/107106;
https://conservancy.umn.edu/handle/11299/107106.

Arden:1960:LPR

[AGG60] B. Arden, B. Galler, and R. Graham. Letters: Preliminary report (February 4, 1960) of Peter Naur. Communications of the ACM, 3(6):A13, June 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [NBB^{+}60].

Alefeld:2006:BRJ

[Ale06] G. Alefeld. Book review: Josef Stoer und Roland Bulirsch, Numerische Mathematik 2. Eine Einführung — unter Berücksichtigung von Vorlesungen von F. L. Bauer. Zeitschrift für Angewandte Mathematik und Mechanik, 86(7):584, July 2006. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Anonymous:2004:FB

[Ano04] Anonymous. F. L. Bauer 80. Informatik Spektrum, 27(4):??, ???? 2004. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-004-0414-9.

Anonymous:2005:CPSa

[Ano05] Anonymous. Call for papers: Special issue in honor of Friedrich L. Bauer. Linear Algebra and its Applications, 394(1):351, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2016:FB

[Ano16] Anonymous. Friedrich L. Bauer. Audio presentation of 9m29s on F. L. Bauer., May 24, 2016.

Anonymous:2019:FBB

[Ano19] Anonymous. Friedrich L. Bauer: biography. Web site, May 18,

2019. URL https://fampeople.com/cat-friedrich-l-bauer.

Bauer:1949:AWE

[Bau49a] F. L. Bauer. Ausreduzierte Wellengleichungen für Elementarteilchen von halbzahligem Spin. (German) [Reduced wave equations for elementary particles of half-integer spin]. Zeitschrift für Naturforschung, Teil A: Astrophysik, Physik und Physikalische Chemie, 4(??):720–721, ???? 1949. CODEN ZENAAU. ISSN 0044-3166.

Bauer:1949:ALG

[Bau49b] F. L. Bauer. Zur Algebraik des Logikkalküls. (German) [On the algebra of logic calculus]. Methodos, 1:288–292, 1949. CODEN MTHDA5. ISSN 0369-2361.

Bauer:1950:MPG

[Bau50] F. L. Bauer. Magnetische Punktpole. (German) [Magnetic point poles]. Physik Journal, 6(6):257–261, June 1950. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Bauer:1951:SAG

[Bau51] F. L. Bauer. Das Schalenmodell der Atomkerne. (German) [The shell model of the atomic nucleus]. Physik Journal, 7(5):205–214, May 1951. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Bauer:1952:GUTb

[Bau52a] Friedrich L. Bauer. Gruppentheoretische Untersuchungen zur Theorie der Spinwellengleichungen. (German) [Group-theoretical studies on the theory of spin wave equations]. Sitzungsberichte der Mathematisch-Naturwissenschaftlichen Klasse der Bayerischen Akademie der Wissenschaften zu München, 1952: 111–179 (1953), 1952. CODEN SMNBAC. ISSN 0376-1037. Author’s Habilitation dissertation.

Bauer:1952:RSF

[Bau52b] Friedrich L. Bauer. Sur les représentations spinorielles. (French) [On spinor representations]. Comptes rendus de l’Académie des sciences, Paris, 234:1743–1744, 1952. ISSN 0001-4036 (print), 2419-6304 (electronic).

Bauer:1952:TDE

[Bau52c] Friedrich L. Bauer. Tenseurs, dont les éléments sont des matrices de Dirac. (French) [Tensors, whose elements are Dirac matrices]. Comptes rendus de l’Académie des sciences, Paris, 235:793–794, 1952. ISSN 0001-4036 (print), 2419-6304 (electronic).

Bauer:1952:GUTa

[Bau52d] Friedrich Ludwig Bauer. Gruppentheoretische Untersuchungen zur Theorie der Spinwellengleichungen. (German) [Group theoretical studies on the theory of spin wave equations]. Ph.D. (Dr. rer. nat.) dissertation, Ludwig-Maximilians-Universität, Munich, West Germany, 1952. Published in [Bau52a].

Bauer:1953:PST

[Bau53] Friedrich L. Bauer. Process for securing the transmission of messages pulse groups to disturbances. German patent DE892767C.,

October 8, 1953. URL https://patents.google.com/patent/

DE892767C. Patent filed 20 January 1951.

Bauer:1954:QKB

[Bau54a] F. L. Bauer. Quadratische Konvergenz der Bernoullischen Methode. (German) [Quadratic convergence of Bernoulli’s method]. Zeitschrift für Angewandte Mathematik und Mechanik, 34(8–9): 283, 1954. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Bauer:1954:BEN

[Bau54b] Friedrich L. Bauer. Beiträge zur Entwicklung numerischer Verfahren für programmgesteuerte Rechenanlagen. I. Quadratisch konvergente Durchführung der Bernoulli-Jacobischen Methode zur Nullstellenbestimmung von Polynomen. (German) [Contributions to the development of numerical methods for programcontrolled computing systems. I. Quadratically convergent execution of the Bernoulli–Jacobian method for determining the roots of polynomials]. Sitzungsberichte der Mathematisch-Naturwissenschaftlichen Klasse der Bayerischen Akademie der Wissenschaften zu München, 1954:275–303 (1955), 1954. CO-DEN SMNBAC. ISSN 0376-1037.

Bauer:1954:RAM

[Bau54c] Friedrich L. Bauer. Le rang de l’algébre de matrices se trouvant dans la méthode de fusion de M. Louis de Broglie. (French) [The rank of the matrix algebra in the fusion method of M. Louis de Broglie]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 238(??):1477–1478, ???? 1954. CO-DEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Bauer:1954:PST

[Bau54d] Friedrich L. Bauer. Process for securing the transmission of messages pulse groups. German patent DE1953S0032347., Novem-

ber 22, 1954. URL https://patents.google.com/patent/

DE920368C/. Patent filed 26 February 1953.

Bauer:1954:TSG

[Bau54e] Friedrich L. Bauer. Zur Theorie der Spingruppen. (German) [On the theory of the spin groups]. Mathematische Annalen, 128: 228–256, 1954. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Bauer:1954:BQK

[Bau54f] Friedrich Ludwig Bauer. Über quadratisch konvergente Iterationsverfahren zur Lösung von algebraischen Gleichungen und Eigenwertproblemen. (German) [On quadratically convergent iteration methods for solving algebraic equations and eigenvalue problems]. Habilitation thesis, Technical Universit of Munich (TUM), Munich, West Germany, 1954.

Bauer:1955:NPQ

[Bau55a] Friedrich L. Bauer. Der Newton-Prozess als quadratisch konvergente Abkürzung des allgemeinen linearen stationären Iterationsverfahrens 1. Ordnung (Wittmeyer-Prozess). (German) [The Newton process as a quadratic convergent abbreviation abbreviation of the general linear stationary iteration process of first order (Wittmeyer process)]. Zeitschrift für Angewandte Mathematik und Mechanik, 35(12):469–470, 1955. CODEN ZA-MMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Bauer:1955:BKV

[Bau55b] Friedrich L. Bauer. Die Betriebs-Kettenmatrix von Vierpolen. (German) [The operational chain matrix of quadrupoles]. Arch. Elek. Übertr., 9:559–560, 1955. ISSN 0374-2393.

Bauer:1955:DIH

[Bau55c] Friedrich L. Bauer. Ein direktes Iterationsverfahren zur Hurwitz-Zerlegung eines Polynoms. (German) [A direct iteration method for the Hurwitz decomposition of a polynomial]. Arch. Elek. Übertr., 9:285–290, 1955. ISSN 0374-2393.

Bauer:1955:PST

[Bau55d] Friedrich L. Bauer. Process for securing the transmission of messages pulse groups. German patent DE924571C., March 3,

1955. URL https://patents.google.com/patent/DE924571C.

Patent filed 16 July 1953.

Bauer:1956:NBA

[Bau56a] F. L. Bauer. Zur numerischen Behandlung von algebraischen Eigenwertproblemen höherer Ordnung. (German) [On the numerical treatment of algebraic eigenvalue problems of higher order]. Zeitschrift für Angewandte Mathematik und Mechanik, 36(7–8):244–245, 1956. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Bauer:1956:BID

[Bau56b] Friedrich L. Bauer. Beiträge II, Direkte Faktorisierung eines Polynoms. (German) [Contributions ii, direct factoring of a polynomial]. Sitzungsberichte — Bayerische Akademie der Wissenschaften, Mathematisch-Naturwissenshaftliche Klasse, 1956: 163–203, 1956. CODEN SMNBAC. ISSN 0340-1037.

Bauer:1956:BEN

[Bau56c] Friedrich L. Bauer. Beiträge zur Entwicklung numerischer Verfahren für programmgesteuerte Rechenanlagen. II. Direkte Faktorisierung eines Polynoms. (German) [Contributions to the development of numerical methods for program-controlled computing systems. II. Direct factorization of a polynomial]. Bayer.

Akad. Wiss. Math.-Nat. Kl. S.-B., 1956:163–203 (1957), 1956. ISSN 0340-7586.

Bauer:1956:VAI

[Bau56d] Friedrich L. Bauer. Das Verfahren der abgekürzten Iteration für algebraische Eigenwertprobleme, insbesondere zur Nullstellenbestimmung eines Polynoms. (German) [The method of the abbreviated iteration for algebraic eigenvalue problems, in particular for determining the roots of a polynomial]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 7:17–32, 1956. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Bauer:1957:BDV

[Bau57a] F. L. Bauer. Beiträge zum Danilewski-Verfahren. (German) [Contributions to the Danilewski process]. In Aktuelle Problem der Rechentechnik. Bericht über das Internationale Mathematiker-Kolloquium, Dresden, 22. bis 27. November 1955. (German) [Current problem of computing technology. Report on the International Mathematician Colloquium, Dresden, 22nd to 27th November 1955], pages 133–139. VEB Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1957.

Bauer:1957:ZZE

[Bau57b] F. L. Bauer. Zusammenhänge zwischen einigen Iterationsverfahren der linearen Algebra. Aktuelle Probleme der Rechentechnik. (German) [Relationships between some iteration methods of linear algebra. Current problems of computing technology]. In Bericht über das Internationale Mathematiker-Kolloquium, Dresden, 22. bis 27. November 1955, pages 99–111. VEB Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1957.

Bauer:1957:VTV

[Bau57c] Friedrich L. Bauer. Das Verfahren der Treppeniteration und verwandte Verfahren zur Lösung algebraischer Eigenwertprobleme. (German) [The method of step iteration and related methods for solving algebraic eigenvalue problems]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 8:214–235, 1957. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Bauer:1958:MMI

[Bau58a] F. L. Bauer. On modern matrix iteration processes of Bernoulli and Graeffe type. Journal of the ACM, 5(3):246–257, July 1958.

CODEN JACOAH. ISSN 0004-5411 (print), 1557-735x (electronic).

Bauer:1958:MRI

[Bau58b] Friedrich L. Bauer. Method for removing interference telegraphic messages. Swiss patent CH327525A., January 31,

1958. URL https://patents.google.com/patent/CH327525A.

Patent filed 21 December 1954.

Bauer:1959:SRT

[Bau59a] F. L. Bauer. Sequential reduction to tridiagonal form. Journal of the Society for Industrial and Applied Mathematics, 7(1):107– 113, March 1959. CODEN JSIMAV. ISSN 0368-4245 (print),

1095-712X (electronic). URL https://www.jstor.org/stable/
2099007.

Bauer:1959:MAS

[Bau59b] Friedrich L. Bauer. Method of and apparatus for securing the transmission of telegraph impulses. US patent US2880273A., March 31, 1959. Patent filed 12 July 1954.

Bauer:1959:QDE

[Bau59c] Friedrich L. Bauer. The quotient–difference and epsilon algorithms. In Langer [Lan59], pages x + 462. LCCN QA3 .U45 no.

1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Bauer:1960:FCL

[Bau60a] F. L. Bauer. The formula-controlled logical computer “Stanislaus”. Mathematics of Computation, 14(69):64–67, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.jstor.org/stable/2002986. Tech-

nical Notes and Short Papers.

Bauer:1960:A

[Bau60b] Friedrich L. Bauer. The g-algorithm. Journal of the Society for Industrial and Applied Mathematics, 8(1):1–17, March 1960. CO-DEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

URL https://www.jstor.org/stable/2098950.

Bauer:1961:ARI

[Bau61a] F. L. Bauer. Algorithm 60: Romberg integration. Communications of the ACM, 4(6):255, June 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1961:FGK

[Bau61b] F. L. Bauer. A further generalization of the Kantorovič inequality. Numerische Mathematik, 3:117–119, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1961:IDC

[Bau61c] F. L. Bauer. Influence du choix du pivot sur l’erreur d’arrondi. (French) [Influence of the choice of pivot on the rounding error]. In ????, editor, 1er Congrè de l’Association Française de Calcul AFCAL, Grenoble, 14–15–16 septembre 1960, pages 145– 150. ????, ????, 1961.

Bauer:1961:MIN

[Bau61d] F. L. Bauer. La méthode d’intégration numérique de Romberg. (French) [Romberg numerical integration method]. In Centre Belge de Recherche en Mathématiques. Colloque Analyse Numérique, (Mons, 1961), pages 119–129. Librairie Universitaire, Louvain, Belgium, 1961.

Bauer:1961:WHN

[Bau61e] F. L. Bauer. Wertevorrat eines Hauptabschnitts einer normalen Matrix. (German) [Stock of [eigen]values of a main section of a normal matrix]. Jber. Deutsch. Math.–Verein, 63(Abt. 1):180– 182 (1961), 1961. ISSN 0012-0456 (print), 1869-7135 (electronic).

Bauer:1962:FVS

[Bau62] F. L. Bauer. On the field of values subordinate to a norm. Numerische Mathematik, 4:103–113, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1963:AG

[Bau63a] F. L. Bauer. Algorithm 153: Gomory. Communications of the ACM, 6(2):68, February 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1963:OSMb

[Bau63b] F. L. Bauer. Optimal scaling of matrices and the importance of the minimal condition. In Cicely M. Popplewell, editor, Proceedings of the IFIP Congress Munich 27 August–1 September 1962, Information Processing 62, pages 198–201. North-Holland, Amsterdam, The Netherlands, Amsterdam, The Netherlands, 1963.

Bauer:1963:OSMa

[Bau63c] F. L. Bauer. Optimally scaled matrices. Numerische Mathematik, 5(1):73–87, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1964:RCP

[Bau64] F. L. Bauer. Reply to comments by P. A. Samet and P. J. Taylor. Communications of the ACM, 7(7):447, July 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1965:BLP

[Bau65a] F. L. Bauer. Buchbesprechung: A. A. Ljapunow, Probleme der Kybernetik, Band 1. 326 S. m. 23 Abb. Berlin 1962. Akademie-Verlag. Preis geb. MDN 44,–. Zeitschrift für Angewandte Mathematik und Mechanik, 45(1):68, 1965. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Bauer:1965:EPH

[Bau65b] F. L. Bauer. An elementary proof of the Hopf inequality for positive operators. Numerische Mathematik, 7(4):331–337, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1965:EWR

[Bau65c] F. L. Bauer. Elimination with weighted row combinations for solving linear equations and least squares problems. Numerische Mathematik, 7(4):338–352, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https:/

/link.springer.com/chapter/10.1007/978-3-642-86940-2_
9;   https://link.springer.com/chapter/10.1007/978-3-

662-39778-7_9. Also in [WR71, pp. 119–133].

Bauer:1965:NST

[Bau65d] F. L. Bauer. Nonlinear sequence transformations. In Approximation of Functions (Proc. Sympos. General Motors Res. Lab., 1964 ), pages 134–151. Elsevier, Amsterdam, The Netherlands, 1965.

Bauer:1965:NAB

[Bau65e] F. L. Bauer. Numerische Abschätzung und Berechnung von Eigenwerten nichtsymmetrischer Matrizen. (German) [Numerical estimation and calculation of eigenvalues non-symmetrical matrices]. Aplikace Matematiky, 10:178–189, 1965. CODEN APM-TAK. ISSN 0373-6725.

Bauer:1965:UCF

[Bau65f] F. L. Bauer. Use of continued fractions and algorithms related to them. Quest. Analisi numer., ??(??):351–419, 1965.

Bauer:1966:GLL

[Bau66a] F. L. Bauer. Genauigkeitsfragen bei der Lösung linearer Gleichungssysteme. (German) [Questions of accuracy in solving linear systems of equations]. Zeitschrift für Angewandte Mathematik und Mechanik, 46(7):409–421, November 1966. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/10.
1002/zamm.19660460702.

Bauer:1966:ISR

[Bau66b] F. L. Bauer. Inclusion sets for roots of polynomials. Numerische Mathematik, 9(2):173–176, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1966:SSF

[Bau66c] F. L. Bauer. Syntaktische Struktur formaler und natürlicher Sprachen. (German) [Syntactic structure formulas and natural languages]. Computing, 1(2):93–104, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

https://link.springer.com/article/10.1007/BF02342618.

Bauer:1968:FVG

[Bau68a] F. L. Bauer. Fields of values and Gershgorin disks. Numerische Mathematik, 12(2):91–95, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1968:SAS

[Bau68b] F. L. Bauer. Some aspects of scaling invariance. In Programmation en Mathématiques Numériques (Actes Colloq. Internat. C.N.R.S. No. 165, Besançon, 1966), pages 37–47. Éditions du Centre National de Recherche Scientifique, Paris, France, 1968.

Bauer:1968:VNM

[Bau68c] Friedrich L. Bauer. Vorträge über numerische Mathematik. (German) [Lectures on numerical mathematics]. Wissenschaftliche Zeitschrift der Technischen Universität Dresden, 17:1127–1129, 1968. CODEN WZTUAU. ISSN 0043-6925.

Bauer:1969:ROS

[Bau69] F. L. Bauer. Remarks on optimally scaled matrices. Numerische Mathematik, 13(1):1–3, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1971:HR

[Bau71] F. L. Bauer. Heinz Rutishauser (10.11.1970). Computing, 7(1– 2):129–130, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL https://link.springer.

com/article/10.1007/BF02279949.

Bauer:1972:SE

[Bau72] Friedrich L. Bauer. Software engineering. In Freiman et al. [FGR72], pages 530–538. ISBN 0-7204-2063-6. LCCN QA75.5

.I532 1971. URL https://dblp.org/db/conf/ifip/ifip71-1.
html.

Bauer:1973:ACS

[Bau73a] F. L. Bauer. Advanced Course on Software Engineering, volume 81 of Lecture Notes in Economics and Mathematical Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. ISBN 0-387-06185-1, 3-540-06185-1. xii + 545 pp. LCCN QA76.6 .A33 1972.

Bauer:1973:SSE

[Bau73b] F. L. Bauer. Software and software engineering. SIAM Review, 15(2):469–480, April 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://www.jstor.org/

stable/2028680.

Bauer:1974:AHR

[Bau74a] F. L. Bauer. Appendix: Historical remarks on compiler construction. In Compiler Construction: an Advanced Course [BDG^{+}74], pages 603–621. ISBN 3-540-06958-5 (print), 3-662-21549-7 (ebook). LCCN QA76.6 .C6281. URL https://link.springer.

com/book/10.1007/978-3-662-21549-4.

Bauer:1974:CGR

[Bau74b] F. L. Bauer. Computational graphs and rounding error. SIAM Journal on Numerical Analysis, 11(1):87–96, March 1974. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

URL https://www.jstor.org/stable/2156433. In memoriam

George Forsythe.

Bauer:1974:WHW

[Bau74c] F. L. Bauer. Was heißt und was ist Informatik?. (German) [What does computer science mean, and what is it?]. In Mathematiker über Math., pages 349–368. ????, ????, 1974.

Bauer:1974:GVS

[Bau74d] Friedrich L. Bauer, editor. Gatlinburg VI: Symposium on Numerical Algebra, December 15–22, 1974, Hopfen am See, near Munich, Germany. Abstracts. Technische Universität, Institut für Mathematik, Munich, West Germany, 1974. LCCN ????

Bauer:1975:PN

[Bau75] Friedrich L. Bauer. Positivity and norms. Communications of the ACM, 18(1):9–13, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder.

Bauer:1976:PP

[Bau76a] F. L. Bauer. A philosophy of programming. In Bauer and Samelson [BS76], pages 194–241. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L335. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0046.htm;   http://www.springerlink.com/content/978-
0-387-07994-3;   http://www.springerlink.com/openurl.
asp?genre=book&isbn=978-3-540-07994-1;   http://www.
springerlink.com/openurl.asp?genre=issue&issn=0302-9743&
volume=46.

Bauer:1976:PPC

[Bau76b] F. L. Bauer. A philosophy of programming: A course of 3 lectures given at the Imperial College of Science and Technology, University of London on 3, 4 and 5 October 1973. In Bauer and Samelson [BS76], pages 194–241. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.L335. URL https://link.springer.com/chapter/10.1007/
3-540-07994-7_55.

Bauer:1976:PEPa

[Bau76c] F. L. Bauer. Programming as an evolutionary process. In Bauer and Samelson [BS76], pages 153–182. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.6 .L335. URL http://link.springer-ny.
com/link/service/series/0558/tocs/t0046.htm;   http:/
/www.springerlink.com/content/978-0-387-07994-3;   http:
//www.springerlink.com/openurl.asp?genre=book&isbn=978-
3-540-07994-1;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=46.

Bauer:1976:PEPb

[Bau76d] F. L. Bauer. Programming as an evolutionary process. In Proceedings of the 2nd International Conference on Software Engineering, San Francisco, 1976, pages 223–234. ????, ????, 1976.

Bauer:1976:WPG

[Bau76e] F. L. Bauer. Was ist Programmtransformation?. (German) [What is program transformation?]. Elektronische Rechenanlagen, 18(??):229–233, ???? 1976. CODEN ELRAA4. ISSN 0013-5720.

Bauer:1977:AMC

[Bau77a] F. L. Bauer. Angstl’s mechanism for checking wellformedness of parenthesis-free formulae. Mathematics of Computation, 31 (137):318–320, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.jstor.

org/stable/2005801.

Bauer:1977:AHR

[Bau77b] F. L. Bauer. Appendix: Historical remarks on compiler construction. In Compiler construction: an advanced course [BE76], page ?? CODEN LNCSD9. ISBN 0-387-08046-5 (USA), 3-540-08046-5 (Berlin), 0-387-07542-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .C628 1976. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0021.htm;   http://www.springerlink.com/content/978-
0-387-08046-8;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=21.

Bauer:1977:SEA

[Bau77c] Friedrich Ludwig Bauer, editor. Software Engineering: an Advanced Course. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second [reprint of first] edition, 1977. ISBN 0-387-08364-2, 3-540-08364-2. xii + 545 pp. LCCN QA76.6 .A33 1972.

Bauer:1978:DPL

[Bau78] F. L. Bauer. Design of a programming language for a program transformation system. In Sigram Schindler and Wolfgang K. Giloi, editors, GI — 8. Jahrestagung, Berlin, volume 16, pages 1–28. Springer-Verlag, Berlin, Germany / Heidelberg, Germany /

London, UK / etc., 1978. URL https://link.springer.com/
chapter/10.1007/978-3-642-67091-6_1.

Bauer:1979:DLE

[Bau79a] F. L. Bauer. Detailization and lazy evaluation, infinite objects and pointer representation. In Bauer and Broy [BB79], pages 406–420. CODEN LNCSD9. ISBN 0-387-09251-X (print), 3-540-35312-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0069.htm;   http://www.springerlink.com/content/978-
0-387-09251-5;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=69;   https://
link.springer.com/book/10.1007/BFb0014649. The interna-

tional summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.

Bauer:1979:SIT

[Bau79b] F. L. Bauer. From specification to implementation — the formal approach. In Bauer and Broy [BB79], pages 235– 236. CODEN LNCSD9. ISBN 0-387-09251-X (print), 3-540-35312-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http://link.

springer-ny.com/link/service/series/0558/tocs/t0069.
htm;   http://www.springerlink.com/content/978-0-387-
09251-5;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=69;   https://link.springer.
com/book/10.1007/BFb0014649. The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.

Bauer:1979:PDS

[Bau79c] F. L. Bauer. Program development by stepwise transformations — the project CIP. In Bauer and Broy [BB79], pages 237–272. CODEN LNCSD9. ISBN 0-387-09251-X (print), 3-540-35312-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0069.htm;   http://www.springerlink.com/content/978-
0-387-09251-5;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=69;   https://
link.springer.com/book/10.1007/BFb0014649. The interna-

tional summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.

Bauer:1980:BZR

[Bau80] Friedrich L. Bauer. Between Zuse and Rutishauser — the early development of digital computing in Central Europe. In Metropolis et al. [MHR80], pages 505–524. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. URL http://

www.sciencedirect.com/science/article/pii/B978012491650050037X.

Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Bauer:1981:AA

[Bau81] F. L. Bauer. Algorithms and algebra. In Ershov and Knuth [EK81], pages 421–429. CODEN LNCSD9. ISBN 0-387-11157-3 (New York), 3-540-11157-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .A43 1981.

URL https://link.springer.com/chapter/10.1007/3-540-

11157-3_39. The symposium was organized by the Academy of Sciences of the Uzbek S.S.R.

Bauer:1982:AAR

[Bau82a] F. L. Bauer. Algorithms and algebra. (Russian). In ????, editor, Algorithms in modern mathematics. II Mater. Int. Symp. Urgench/UzSSR, 1979 No. 2, pages 230–239. ????, ????, ????

Bauer:1982:KVM

[Bau82b] Friedrich L. Bauer. Kryptologie — Verfahren und Maximen. (German) [Cryptology — procedures and maxims]. Informatik Spektrum, 5(??):74–81, ???? 1982. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Bauer:1985:KZJ

[Bau85a] F. L. Bauer. Konrad Zuse 75 Jahre. (German) [Konrad Zuse 75 years old]. Elektronische Rechenanlagen, 27(??):129–131, 1985. CODEN ELRAA4. ISSN 0013-5720.

Bauer:1985:WAD

[Bau85b] F. L. Bauer. Warum abstrakte Datentypen?. (German) [Why abstract data types?]. Informatik Spektrum, pages 29–36, ???? 1985. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Bauer:1987:LC

[Bau87] F. L. Bauer. Logic and computers. In Logik, Wissenschaftstheorie und Erkenntnistheorie (Kirchberg am Wechsel, 1986), volume 13 of Schriftenreihe Wittgenstein-Ges., pages 60–67. Hölder-Pichler-Tempsky, Vienna, Austria, 1987.

Bauer:1988:APG

[Bau88a] Friedrich L. Bauer. For all primes greater than 3, ({2p−1 \atop p−1}) ≡ 1 (mod p^{3}) holds. The Mathematical Intelligencer, 10(3):42, 1988. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Bauer:1988:ZJI

[Bau88b] Friedrich L. Bauer. Zwanzig Jahre Institut für Informatik: Informatik und Algebra. (German) [Twenty years of the Institute for Informatics: Informatics and algebra]. Report, ETH Zürich, Zürich, Switzerland, October 16, 1988. 7–

24 pp. URL http://pascal.hansotten.com/uploads/wirth/
Zwanzig%20Jahre%20Institut%20Informatik%20095.pdf.

Bauer:1989:BKR

[Bau89a] Friedrich L. Bauer. Eine Bemerkung zu Koechers Reihen für die Eulersche Konstante. (German) [A remark on Koecher’s series for the Euler’s constant]. Bayer. Akad. Wiss. Math.-Natur. Kl. Sitzungsber., pages 27–33 (1990), 1989. ISSN 0340-7586.

Bauer:1989:MAP

[Bau89b] Friedrich L. Bauer. In memoriam: Andrei Petrovich Ershov, 19 April 1931–8 December 1988. Acta Informatica, 27(3): 193–194, December 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/BF00572987.

Bauer:1990:CPS

[Bau90a] F. L. Bauer. The cellar principle of state transition and storage allocation. Annals of the History of Computing, 12(1): 41–49, January/March 1990. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1990/
pdf/a1041.pdf;   http://www.computer.org/annals/an1990/
a1041abs.htm.

Bauer:1990:ESN

[Bau90b] F. L. Bauer. Efficient solution of a non-monotonic inverse problem. In Feijen et al. [FvGGM90], chapter 2, pages 19– 26. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL https://link.springer.com/chapter/10.1007/

978-1-4612-4476-9_3. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Bauer:1990:KAG

[Bau90c] F. L. Bauer. Kurzer Abrißder Geschichte der Informatik 1890–1990. (German) [Brief outline of the history of IT 1890– 1990]. In Gerd Fischer, editor, Ein Jahrhundert Mathematik 1890–1990. Festschrift zum Jubiläum der DMV, volume 6 of Dokumente Gesch. Math., pages 113–147. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1990. URL https:/

/link.springer.com/chapter/10.1007/978-3-322-80265-1_

3. Deutsche Mathematiker-Vereinigung e.V. Freiburg im Breisgau (DE).

Bauer:1991:LAC

[Bau91a] Friedrich L. Bauer, editor. Logic, algebra, and computation. Proceedings, Marktoberdorf, FRG, July 25–August 6, 1989, volume 79 of NATO ASI Series F. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-642-76801-6 (print), 3-642-76799-0 (e-book). LCCN QA76.9.L63.

Bauer:1991:SH

[Bau91b] Friedrich L. Bauer. Sternpolygone und Hyperwürfel. (German) [Star polygons and hypercubes]. In Miscellanea mathematica, pages 7–44. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Bauer:1992:SMD

[Bau92] Friedrich Ludwig Bauer. Strukturerkennung in der Mathematik und durch die Mathematik. (German) [Structure recognition in mathematics and through mathematics]. Nova Acta Leopoldina (N.F.), 67(281):97–119, 1992. CODEN NOALA4. ISSN 0369-5034.

Bauer:1993:KMM

[Bau93] Friedrich L. Bauer. Kryptologie: Methoden und Maximen. (German) [Cryptology: methods and maxims]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56356-3 (print), 3-540-57771-8 (print), 3-642-78869-6, 3-642-97476-7. xii + 357 pp. LCCN QA76.9.A25.

Bauer:1994:MDS

[Bau94a] F. L. Bauer. Multiplication and dual system. Informatik Spektrum, 17(4):245–250, August 1994. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Bauer:1994:KMM

[Bau94b] Friedrich L. Bauer. Kryptologie: Methoden und Maximen. (German) [Cryptology: methods and maxims]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1994. ISBN 3-642-78869-6, 3-642-97476-7. xii +

357 pp. LCCN QA76.9.A25. URL http://www.springerlink.
com/content/978-3-642-78869-7.

Bauer:1995:BRC

[Bau95a] F. L. Bauer. Book reviews: Cap, C. H., 310447Theoretische Grundlagen der Informatik. Wien/New York, Springer-Verlag 1993. xiii + 332 S. ISBN 3-211-82527-4. Zeitschrift für Angewandte Mathematik und Mechanik, 75(7):562, 1995. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Bauer:1995:LES

[Bau95b] F. L. Bauer. Letter to the editor: a series from the Karan.apaddhati. The Mathematical Intelligencer, 17(3):3–4, January 1995. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Bauer:1995:EGM

[Bau95c] Friedrich L. Bauer. Entzifferte Geheimnisse: Methoden und Maximen der Kryptologie. (German) [Deciphering secrets: methods and maxima of cryptology]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58118-9, 3-540-62632-8, 3-642-97579-8, 3-642-97580-1. xii + 391 pp. LCCN Z103 .B378 1997. URL http://www.gbv.de/

dms/goettingen/16610728X.pdf.

Bauer:1996:CDM

[Bau96a] Friedrich L. Bauer. Cryptological devices and machines in the Deutsches Museum, Munich. Cryptologia, 20(1):11–13, 1996. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Bauer:1996:KZF

[Bau96b] Friedrich L. Bauer. Konrad Zuse: Fakten und Legenden. (German) [Konrad Zuse: facts and legends]. Informatik Spektrum, 19(6):??, ???? 1996. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s002870050040.

Bauer:1996:KBG

[Bau96c] Friedrich L. Bauer. Kryptologie und Blindenschrift. (German) [Cryptology and Braille]. Informatik Spektrum, 19(1):??, ???? 1996. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X

(electronic). URL http://link.springer.com/article/10.
1007/s002870050015.

Bauer:1996:PKG

[Bau96d] Friedrich L. Bauer. Punkt und Komma. (German) [Period and comma]. Informatik Spektrum, 19(2):93–95, April 1996. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s002870050023.

Bauer:1996:ZGM

[Bau96e] Friedrich L. Bauer. Zaubergemurmel. (German) [Magic murmur]. Informatik Spektrum, 19(5):??, ???? 1996. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://

link.springer.com/article/10.1007/s002870050039.

Bauer:1997:MFT

[Bau97a] F. L. Bauer. The might of formulas and their limits. Lecture Notes in Computer Science, 1337:1–8, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.com/link/service/series/
0558/bibs/1337/13370001.htm;   http://link.springer-
ny.com/link/service/series/0558/papers/1337/13370001.
pdf;   https://link.springer.com/chapter/10.1007/BFb0052072.

Bauer:1997:DSM

[Bau97b] Friedrich L. Bauer. Decrypted Secrets: Methods and Maxims of Cryptology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-60418-9. xii + 447 pp. LCCN QA76.9.A25 B38513 1997. Methods and maxims of cryptology.

Bauer:1997:EGM

[Bau97c] Friedrich L. Bauer. Entzifferte Geheimnisse: Methoden und Maximen der Kryptologie. (German) [Deciphered secrets: methods and maxims of cryptology]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second expanded edition, 1997. ISBN 3-642-97972-6, 3-642-97973-4 (print). xiv + 472 pp. LCCN QA76.9.A25.

Bauer:1997:GAL

[Bau97d] Friedrich L. Bauer. Genesis of algorithmic languages. In Mathematical methods in program development. Proceedings of the NATO ASI, Marktoberdorf, Germany, July 30–August 11 1996, volume 158 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 215–269. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. With an appendix in German by K. Samelson.

Bauer:1997:PLH

[Bau97e] Friedrich L. Bauer. Pringsheim, Liebmann, Hartogs—Schicksale jüdischer Mathematiker in München. (German) [Pringsheim, Liebmann, Hartogs — fates of Jewish mathematicians in Munich]. Bayer. Akad. Wiss. Math.-Natur. Kl. Sitzungsber., pages 1–32, 1997. ISSN 0340-7586.

Bauer:1998:OTT

[Bau98a] Friedrich L. Bauer. Olga Taussky-Todd 30.8.1906–7.10.1995. Linear Algebra and its Applications, 280(1):9–12, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Bauer:1998:WEN

[Bau98b] Friedrich L. Bauer. Wer erfand den von-Neumann-Rechner?. (German) [Who invented the von Neumann calculator?]. Informatik Spektrum, 21(2):??, ???? 1998. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s002870050091.

Bauer:1998:ZAE

[Bau98c] Friedrich L. Bauer. Zuse, Aiken und der einschrittige Übertrag. (German) [Zuse, Aiken and the one-step carryover]. Informatik Spektrum, 21(5):??, ???? 1998. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s002870050107.

Bauer:1999:TPG

[Bau99a] Friedrich L. Bauer. Der typographische Punkt. (German) [The typographical point]. Informatik Spektrum, 22(1):41–42, February 1999. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s002870050123.

Bauer:1999:EHR

[Bau99b] Friedrich L. Bauer. An error in the history of rotor encryption devices. Cryptologia, 23(3):206–210, 1999. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Bauer:1999:IIG

[Bau99c] Friedrich L. Bauer. Intuitionismus und Informatik. (German) [Intuitionism and computer science]. Informatik Spektrum, 22(4): ??, ???? 1999. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s002870050145.

Bauer:2000:DS

[Bau00a] Friedrich L. Bauer. Decrypted secrets: Methods and Maxims of Cryptology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., extended edition, 2000. ISBN 3-540-66871-3. xii + 471 pp. LCCN QA76.9.A25. URL http:/

/link.springer.com/10.1007/978-3-662-04024-9. Methods

and maxims of cryptology.

Bauer:2000:DSM

[Bau00b] Friedrich L. Bauer. Decrypted Secrets: Methods and Maxims of Cryptology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second revised and extended edition, 2000. ISBN 3-662-04026-3, 3-662-04024-7 (e-book). xii + 471 pp. LCCN QA76.9.A25. URL http://link.springer.

com/10.1007/978-3-662-04024-9.

Bauer:2000:EGM

[Bau00c] Friedrich L. Bauer. Entzifferte Geheimnisse. Methoden und Maximen der Kryptologie. (German) [Deciphering Secrets: Methods and Maxima of Cryptology]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third revised and expanded edition, 2000. ISBN 3-540-67931-6, 3-642-63545-8 (print), 3-642-58345-8 (e-book). xiii + 502 pp. LCCN

QA76.9.A25. URL http://www.springerlink.com/content/

978-3-642-58345-2. This work refers to [BF60].

Bauer:2000:MRA

[Bau00d] Friedrich L. Bauer. Marian Rejewski und die Alliierten im Angriff gegen die Enigma. (German) [Marian Rejewski and the Allies attacking the Enigma]. Informatik Spektrum, 23(5):??, ???? 2000. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s002870000120.

Bauer:2000:PKZ

[Bau00e] Friedrich L. Bauer. The Plankalkül of Konrad Zuse revisited. In Rojas and Hashagen [RH00], pages 277–293. ISBN 0-262-18197-5 (hardcover), 0-585-35535-5 (electronic). LCCN QA76.17 .F57 2000.

Bauer:2000:GKI

[Bau00f] Friedrich L. Bauer. Zur Geschichte der Kommission für Informatik der Bayerischen Akademie der Wissenschaften (1962– 1999). (German) [On the history of the Commission for Information Science of the Bavarian Academy of Sciences (1962–1999)]. Sitzungsber. Bayer. Akad. Wiss. Math.-Naturwiss. Kl., 1998– 2000:1–26, 2000.

Bauer:2001:CES

[Bau01a] Friedrich L. Bauer. Claude Elwood Shannon 1916–2001. (German) [Claude Elwood Shannon 1916–2001]. Informatik Spektrum, 24(4):228–229, August 2001. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s002870100168.

Bauer:2001:VAV

[Bau01b] Friedrich L. Bauer. Verfahren zur automatischen Verarbeitung von kodierten Daten und Rechenmaschinen zur Ausübung des Verfahrens. (German) [Method for the automatic processing of coded data and calculating machines to carry out the method]. In Broy and Denert [BD01], pages 29–40. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

Bauer:2002:DS

[Bau02a] Friedrich L. Bauer. Decrypted Secrets: Methods and Maxims of Cryptology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third revised and updated edition, 2002. ISBN 3-540-42674-4 (hardcover), 3-662-04736-5 (e-book). xii + 474 pp. LCCN QA76.9.A25 B38513 2001.

Bauer:2002:MYR

[Bau02b] Friedrich L. Bauer. My years with Rutishauser. In ????, editor, LATSIS Symposium 2002: ETH, Zürich, Switzerland, page ?? ????, ????, 2002. URL http://www.cs.umd.

edu/users/oleary/cggg/bauer.pdf;   https://vimeo.com/
308173562/5b0d69f62f.

Bauer:2003:KZH

[Bau03a] Friedrich L. Bauer. Konrad Zuse in Hopferau im Allgäu Z4 und Plankalkül. (German) [Konrad Zuse in Hopferau im Allgäu Z4 and Plankalkül]. Informatik Spektrum, 26(5):??, ???? 2003. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-003-0348-7.

Bauer:2003:MYR

[Bau03b] Friedrich L. Bauer. My year with Rutishauser. DINAL, 11(1–4): 1–13, 2003.

Bauer:2003:QGQ

[Bau03c] Friedrich L. Bauer. QWERTZU. (German) [QWERTY]. Informatik Spektrum, 26(6):??, ???? 2003. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s00287-003-0357-6.

Bauer:2003:WLM

[Bau03d] Friedrich L. Bauer. Why Legendre made a wrong guess about π(x), and how Laguerre’s continued fraction for the logarithmic integral improved it. The Mathematical Intelligencer, 25(3):7– 11, 2003. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Bauer:2004:G

[Bau04a] Friedrich L. Bauer. 3.14159… und 2.71828. . .. (German) [3.14159… and 2.71828…]. Informatik Spektrum, 27(3):??, ???? 2004. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X

(electronic). URL http://link.springer.com/article/10.
1007/s00287-004-0406-9.

Bauer:2004:MQM

[Bau04b] Friedrich L. Bauer. Magische Quadrate und magische Würfel. (German) [Magic squares and magic cubes]. Informatik Spektrum, 27(1):??, ???? 2004. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s00287-003-0351-z.

Bauer:2004:SDG

[Bau04c] Friedrich L. Bauer. Sackgassen und Durchbrüche. (German) [Dead ends and breakthroughs]. Informatik Spektrum, 27(4):??, ???? 2004. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-004-0405-x.

Bauer:2004:TFG

[Bau04d] Friedrich L. Bauer. Theodor Fromme. (German) [Theodor Fromme]. Informatik Spektrum, 27(2):??, ???? 2004. CO-DEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-004-0356-2.

Bauer:2004:WHF

[Bau04e] Friedrich L. Bauer. War Hindenburg ein Feldherr?. (German) [Was Hindenburg a general?]. Informatik Spektrum, 27(5): 467–469, October 2004. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s00287-004-0435-x.

Bauer:2005:MLC

[Bau05a] Friedrich L. Bauer. De Moivre und Lagrange: Cosinus eines rationalen Vielfachen von π. (German) [De Moivre and Lagrange: Cosine of a rational multiple of π]. Informatik Spektrum, 28(2): ??, ???? 2005. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-005-0474-5.

Bauer:2005:LKG

[Bau05b] Friedrich L. Bauer. Lamberts Kettenbruch. (German) [Lambert’s continued fraction]. Informatik Spektrum, 28(4):??, ???? 2005. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-005-0499-9.

Bauer:2005:MBP

[Bau05c] Friedrich L. Bauer. Mathematik besiegte in Polen die unvernünftig gebrauchte ENIGMA Teil 1: Vorarbeiten in Polen Im Gedenken an Marian Rejewski (16. August 1905). (German) [In Poland, mathematics defeated the unreasonably used ENIGMA. Part 1: Preliminary work in Poland in memory of Marian Rejewski (born August 16, 1905)]. Informatik Spektrum, 28(6):??, ???? 2005. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-005-0045-9.

Bauer:2005:PNU

[Bau05d] Friedrich L. Bauer. Polygraphia Nova et Universalis. Informatik Spektrum, 28(3):??, ???? 2005. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s00287-005-0490-5.

Bauer:2005:PTG

[Bau05e] Friedrich L. Bauer. Pythagoräische Tripel. (German) [Pythagorean triple]. Informatik Spektrum, 28(5):??, ???? 2005. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

URL http://link.springer.com/article/10.1007/s00287-
005-0032-1.

Bauer:2005:RH

[Bau05f] Friedrich L. Bauer. Rutishauser, Heinz. In Neue Deutsche Biographie (NDB). (German) [New German Biography], page 305. Duncker & Humblot, Berlin, Germany, 2005. URL https://

daten.digitale-sammlungen.de/0001/bsb00016410/images/
index.html?seite=319.

Bauer:2006:ERA

[Bau06a] F. L. Bauer. An elementary remark on the accuracy of approximations by regular continued fractions. The Mathematical Intelligencer, 28(3):38–43, 2006. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Bauer:2006:FZS

[Bau06b] Friedrich L. Bauer. Frühe Zeugnisse der ,,software“. (German) [Early testimony on “software”]. Informatik Spektrum, 29(6): 433–441, September 2006. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s00287-006-0114-8.

Bauer:2006:SSF

[Bau06c] Friedrich L. Bauer. ,Simple Simon‘: ein früher elektromechanischer Computer. (German) [“Simple Simon”: an early electromechanical computer]. Informatik Spektrum, 29(4):??, ???? 2006. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-006-0093-9.

Bauer:2006:MBP

[Bau06d] Friedrich L. Bauer. Mathematik besiegte in Polen die unvernünftig gebrauchte ENIGMA Teil 1: Vorarbeiten in Polen Im Gedenken an Marian Rejewski (16. August 1905). (German) [In Poland, mathematics defeated the unreasonably used ENIGMA. Part 1: Preliminary work in Poland in memory of Marian Rejewski (born August 16, 1905)]. Informatik Spektrum, 29(1):??, ???? 2006. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-005-0051-y.

Bauer:2006:SBC

[Bau06e] Friedrich L. Bauer. Seit Bombelli und Cataldi: Periodische Kettenbrüche. (German) [Since Bombelli and Cataldi: Periodic continued fractions]. Informatik Spektrum, 29(5):??, ???? 2006. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-006-0104-x.

Bauer:2007:RSF

[Bau07a] F. L. Bauer. Remark on Stirling’s formula and on approximations for the double factorial. The Mathematical Intelligencer, 29(2): 10–14, March 2007. CODEN MAINDC. ISSN 0343-6993 (print),

1866-7414 (electronic). URL https://link.springer.com/
article/10.1007/BF02986199;   https://link.springer.
com/content/pdf/10.1007/BF02986199.pdf.

Bauer:2007:RMB

[Bau07b] Friedrich Bauer. Rotor machines and bombes. In de Leeuw and Bergstra [dLB07], chapter 14, pages 381–446. ISBN 0-444-51608-5 (hardcover). LCCN Z103 .H63 2007. URL http://www.

sciencedirect.com/science/book/9780444516084.

Bauer:2007:ALG

[Bau07c] Friedrich L. Bauer. e^{p}i und π^{e} (Anläßlich des 300. Geburtstags von Leonhard Euler). (German) [e^{π} and π^{e} (on the occasion of Leonhard Euler’s 300th birthday)]. Informatik Spektrum, 30(5): ??, ???? 2007. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-007-0176-2.

Bauer:2007:DSM

[Bau07d] Friedrich L. Bauer. Decrypted Secrets: Methods and Maxims of Cryptology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2007. ISBN 3-540-24502-2 (hardcover), 3-642-06383-7 (paperback), 3-540-48121-4 (e-book). xii + 524 + 16 pp. LCCN QA76.9.A25 B38513 2007. Methods and maxims of cryptology.

Bauer:2007:FJI

[Bau07e] Friedrich L. Bauer. Festschrift 40 Jahre Informatik in München 1967–2007. (German) [Commemoration of 40 years of informatics in Munich 1967–2007)]. Fakultät für Informatik der TUM, Munich, Germany, 2007. xiv + 530 pp. URL https://www.in.

tum.de/die-fakultaet/geschichte/40-jahre-informatik-
muenchen/.

Bauer:2007:FRE

[Bau07f] Friedrich L. Bauer. Fleissner-Raster und der Erzherzog. (German) [Fleissner-Raster and the Archduke]. Informatik Spektrum, 30(1):36–38, January 2007. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s00287-006-0126-4.

Bauer:2007:GLE

[Bau07g] Friedrich L. Bauer. Gregory–Leibniz und Euler: Arcus-Cotangens-Relationen (Anläßlich des 300. Geburtstags von Leonhard Euler). (German) [Gregory–Leibniz and Euler: Arccotangent relations (on the occasion of the 300th birthday of Leonhard Euler)]. Informatik Spektrum, 30(6):??, ???? 2007. CO-DEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-007-0189-x.

Bauer:2007:MJT

[Bau07h] Friedrich L. Bauer. In memoriam John Todd (1911–2007). Numerische Mathematik, 108(1):1–6, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:2007:RHF

[Bau07i] Friedrich L. Bauer. Richard Hamming: Fehlerkorrigierende Codes. (German) [Richard Hamming: Error correcting codes]. Informatik Spektrum, 30(2):??, ???? 2007. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://

link.springer.com/article/10.1007/s00287-006-0135-3.

Bauer:2007:TTF

[Bau07j] Friedrich L. Bauer. Trits and Trytes: Ein früher ternärer Computer in der Sowjetunion. (German) [Trits and trytes: An early ternary computer in the Soviet Union]. Informatik Spektrum, 30(4):279–284, June 2007. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s00287-007-0166-4.

Bauer:2008:UCB

[Bau08a] Friedrich L. Bauer. Der (ungerade) Collatz-Baum. (German) [The (odd) Collatz tree]. Informatik Spektrum, 31(2):??, ???? 2008. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X

(electronic). URL http://link.springer.com/article/10.
1007/s00287-008-0231-7.

Bauer:2008:EHE

[Bau08b] Friedrich L. Bauer. Erich Hüttenhain: Entzifferung 1939–1945. (German) [Erich Hüttenhain: deciphering 1939–1945]. Informatik Spektrum, 31(3):??, ???? 2008. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s00287-008-0242-4.

Bauer:2008:GWI

[Bau08c] Friedrich L. Bauer. Geschachtelte Wurzeln und ihre Elimination. (German) [Nested roots and their elimination]. Informatik Spektrum, 31(1):??, ???? 2008. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s00287-007-0216-y.

Bauer:2008:EAI

[Bau08d] Friedrich L. Bauer. Über ,,Episoden aus den Anfängen der Informatik an der ETH“von A. Speiser. (German) [On “Episodes from the Beginnings of Computer Science at ETH” by A. Speiser]. Informatik Spektrum, 31(6):600–612, October 2008. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-008-0291-8.

Bauer:2008:WAK

[Bau08e] Friedrich L. Bauer. Wallis-artige Kettenprodukte. (German) [Wallis-style chain products]. Informatik Spektrum, 31(4):??, ???? 2008. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-008-0254-0.

Bauer:2009:EG

[Bau09a] Friedrich Bauer. Entzifferte Geheimnisse. (German) [Deciphered secrets]. In Historische Notizen zur Informatik. (German) [Historical notes on computer science] [Bau09d], pages 124–127. ISBN 3-540-85789-3, 3-540-85790-7. LCCN QA76.17 .B38 2009.

Bauer:2009:FHS

[Bau09b] Friedrich Bauer. Fritz Hartogs Schicksal eines jüdischen Mathematikers in München. (German) [Fritz Hartog’s fate as a Jewish mathematician in Munich]. In Historische Notizen zur Informatik. (German) [Historical notes on computer science] [Bau09d], pages 208–216. ISBN 3-540-85789-3, 3-540-85790-7. LCCN QA76.17 .B38 2009.

Bauer:2009:AIB

[Bau09c] Friedrich L. Bauer. After 1960: Informatics begins to take shape. In Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum [Bau10h], chapter 6, pages 113–121. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2009:HNI

[Bau09d] Friedrich L. Bauer. Historische Notizen zur Informatik. (German) [Historical notes on computer science]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-85789-3, 3-540-85790-7. x + 454 pp. LCCN QA76.17 .B38 2009.

Bauer:2009:KPTa

[Bau09e] Friedrich L. Bauer. Kettenbruch-Phänomene: Teil I: Periodische Kettenbrüche. (German) [Continued fraction phenomena: Part I: Periodic continued fractions]. Informatik Spektrum, 32(1):??, ???? 2009. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-008-0309-2.

Bauer:2009:KPTb

[Bau09f] Friedrich L. Bauer. Kettenbruch-Phänomene: Teil II: Wechselwegnahme für transzendente Zahlen. (German) [Continued fraction phenomena: Part II: Removal of change for transcendental numbers]. Informatik Spektrum, 32(2):??, ???? 2009.

CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-009-0330-0.

Bauer:2010:UCF

[Bau10a] F. L. Bauer. Use of continued fractions and algorithms related to them. In Ghizzetti [Ghi10], pages 352–419. CODEN AINFA2. ISBN 3-642-11026-6 (print), 3-642-11027-4 (e-book). ISSN 0001-5903 (print), 1432-0525 (electronic). LCCN QA297

.C56 2010. URL https://link.springer.com/chapter/10.
1007/978-3-642-11027-6_4.

Bauer:2010:ATM

[Bau10b] Friedrich Bauer. After 1890: In thrall to mechanical and electromechanical devices. In Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum [Bau10h], chapter 4, pages 41–71. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2010:AFL

[Bau10c] Friedrich Bauer. After 1935: Formal languages and algorithms, “universal machines” and electronic solutions. In Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum [Bau10h], chapter 5, pages 73–112. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2010:CUS

[Bau10d] Friedrich Bauer. Calculations using symbols. In Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum [Bau10h], chapter 3, pages 23–40. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2010:CIM

[Bau10e] Friedrich Bauer. Conclusion: Informatics and microelectronics are mutually dependent. In Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum [Bau10h], chapter 7, pages 123–127. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2010:NC

[Bau10f] Friedrich Bauer. Numerical calculations. In Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum [Bau10h], chapter 2, pages 7–21. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2010:P

[Bau10g] Friedrich Bauer. Preamble. In Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum [Bau10h], chapter 1, pages 1–6. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2010:OFC

[Bau10h] Friedrich L. Bauer. Origins and Foundations of Computing: In Cooperation with Heinz Nixdorf MuseumsForum. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (ebook). viii + 142 pp. LCCN QA76.17 .B37 2010. With editorial assistance from Norbert Ryska. Translation from the German language edition: “Kurze Geschichte der Informatik” authored by Friedrich L. Bauer. Copyright Wilhelm Fink Verlag, Paderborn, Germany 2007.

Bauer:2011:AVG

[Bau11a] Friedrich L. Bauer. Die ALGOL-Verschwörung. (German) [The ALGOL conspiracy]. Informatik Spektrum, 35(2):141–149, December 2011. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Bauer:2011:MYR

[Bau11b] Friedrich L. Bauer. My years with Rutishauser. Informatik Spektrum, 34(5):514–518, October 2011. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.

springer.com/article/10.1007/s00287-011-0554-7.

Bauer:2011:PRT

[Bau11c] Friedrich L. Bauer. Promotionsrecht und Rektoratsverfassung an der Technischen Hochschule München. (German) [Right to award doctorates and rectorate constitution at the Technical University of Munich]. Informatik Spektrum, 34(2):220–221, February 2011. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-011-0520-4.

Bauer:2012:AVG

[Bau12] Friedrich L. Bauer. Die ALGOL-Verschwörung. (German) [The ALGOL conspiracy]. Informatik Spektrum, 35(2):141–149, December 2012. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-011-0585-0.

Bauer:1998:GKI

[Bau00] Friedrich L. Bauer. Zur Geschichte der Kommission für Informatik der Bayerischen Akademie der Wissenschaften (1962– 1999). (German) [On the history of the Commission for Computer Science of the Bavarian Academy of Sciences (1962–1999)]. Bayer. Akad. Wiss. Math.-Natur. Kl. Sitzungsber., pages 1–26, 1998/00. ISSN 0340-7586.

Bopp:1949:FWE

[BB49] F. Bopp and F. L. Bauer. Feldmechanische Wellengleichungen für Elementarteilchen verschiedenen Spins. (German) [Field mechanical wave equations for elementary particles with different spins]. Zeitschrift für Naturforschung, Teil A: Astrophysik, Physik und Physikalische Chemie, 4(??):611–625, ???? 1949. CODEN ZENAAU. ISSN 0044-3166.

Bauer:1979:PCI

[BB79] Friedrich Ludwig Bauer and M. Broy, editors. Program Construction: International Summer School, volume 69 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNCSD9. ISBN 0-387-09251-X (print), 3-540-35312-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http://link.

springer-ny.com/link/service/series/0558/tocs/t0069.
htm;   http://www.springerlink.com/content/978-0-387-
09251-5;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=69;   https://link.springer.
com/book/10.1007/BFb0014649. The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.

Broy:1984:SAL

[BB84] Manfred Broy and Friedrich L. Bauer. A systematic approach to language constructs for concurrent programs. Science of Computer Programming, 4(2):103–139, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bauer:2003:EWD

[BB03] Friedrich L. Bauer and Manfred Broy. Edsger W. Dijkstra — Acta Informatica and Marktoberdorf. Acta Informatica, 39(3): 141–142, March 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:1985:MPC

[BBB^{+}85] F. L. Bauer, R. Berghammer, M. Broy, W. Dosch, F. Geiselbrechtinger, R. Gnatz, E. Hangel, W. Hesse, B. Krieg-Brückner, A. Laut, T. Matzner, B. Möller, F. Nickl, H. Partsch, P. Pepper, K. Samelson, M. Wirsing, and H. Wössner. The Munich Project CIP. Vol. I, volume 183 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-15187-7. xi + 275 pp. LCCN QA76.6-76.66. The wide spectrum language CIP-L.

Bauer:1981:PWS

[BBD^{+}81] F. L. Bauer, M. Broy, W. Dosch, R. Gnatz, B. Krieg-Brückner, A. Laut, M. Luckmann, T. Matzner, B. Möller, H. Partsch, P. Pepper, K. Samelson, R. Steinbrüggen, M. Wirsing, and H. Wössner. Programming in a wide spectrum language: a collection of examples. Science of Computer Programming, 1(1– 2):73–114, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Buchwald:1958:BBP

[BBG^{+}58] E. Buchwald, F. L. Bauer, H. Gericke, H. Witting, P. Grassmann, W. Döring, G. Dickel, J. Euler, L. Oster, A. Ebinger, and D. Lilienthal. Bücher Bücher: Planck: Physikalische Abhandlungen und Vorträge/Goethe: Schriften Zur Naturwissenschaft/Erckmann: Via Regia/Courant Und Hilbert: Methods Of Mathematical Physics/Blaschke: Griechische Und Anschauliche Geometrie/Ackermann, Verlag: Einführung In Die Wahr Scheinlichkeitsrechnung/Gray: American Institute Of Physics Handbook/Guggenheim: Thermodynamics/Bruhat: Recueil de Problèmes/Roth und Bekker: Kalorimetrische Methoden Zur Bestimmung Chemischer Reaktionswärmen/Kurzes Lehrbuch Der Physikalischen Chemie/Spitzer: Physics Of Fully Ionized Gases/Wirk und Thilo: Niederfrequenz Und Mittel Frequenz Messtechnik/Hönger und Reuber: Rundfunkröhren/ Mutter: Kompendium Der Photographie/Photographie Und Kinematographie. Physik Journal, 14(5):230–237, May 1958. CO-DEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Backus:1960:RAL

[BBG^{+}60] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language Algol 60. Numerische Mathematik, 2(1):106–136, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.

springer.com/article/10.1007/BF01386216.

Backus:1963:RRAa

[BBG^{+}63a] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. The Computer Journal, 5(4):349–367, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/4/349.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_05/Issue_04/050349.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/349.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/350.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/351.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/352.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/353.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/354.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/355.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/356.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/358.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/359.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/360.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/361.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/363.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/364.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/365.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/366.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/367.tif.

Backus:1963:RRAb

[BBG^{+}63b] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, 4:420–453, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/10.1007/BF01386340. Edited by

Peter Naur. Dedicated to the memory of William Turanski.

Backus:1963:BAS

[BBG^{+}63c] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Bericht über die Algorithmische Sprache ALGOL 60 (ALGOL Bulletin Supplement No. 3). Zur Erinnerung an William Turanski. Herausgegeben von P. Naur. (German) [Report on the algorithmic language ALGOL 60 (ALGOL Bulletin supplement no. 3). In memory of William Turanski. Edited by P. Naur]. Elektronische Datenverarbeitung, 2(??):1–17, 1963. ISSN 0374-3012.

Backus:1976:MRA

[BBG^{+}76] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified report on the algorithmic language ALGOL-60. The Computer Journal, 19(4):364–379, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/364.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/364.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/365.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/366.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/367.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/368.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/369.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/370.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/371.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/372.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/373.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/374.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/375.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/376.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/378.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/379.
tif.

Bauer:1978:TWSa

[BBG^{+}78a] F. L. Bauer, M. Broy, R. Gnatz, W. Hesse, and B. Krieg-Brückner. Towards a wide spectrum language to support program specification and program development. In Klaus Alber, editor, Programmiersprachen: 5. Fachtagung der GI, Braunschweig, 8./9. März 1978, volume 12 of Informatik – Fachberichte, pages 73–85. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL https://link.springer.com/chapter/10.1007/978-3-
642-87956-2_5.

Bauer:1978:TWSb

[BBG^{+}78b] F. L. Bauer, M. Broy, R. Gnatz, W. Hesse, B. Krieg-Brückner, H. Partsch, P. Pepper, and H. Wössner. Towards a wide spectrum language to support program specification and program development. ACM SIGPLAN Notices, 13(12):15–24, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bauer:1979:TWS

[BBG^{+}79] F. L. Bauer, M. Broy, R. Gnatz, W. Hesse, B. Krieg-Brückner, H. Partsch, P. Pepper, and H. Wössner. Towards a wide spectrum language to support program specification and program development. In Bauer and Broy [BB79], pages 543–552. CO-DEN LNCSD9. ISBN 0-387-09251-X (print), 3-540-35312-7 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .P75117. URL https://link.springer.com/chapter/

10.1007/BFb0014682. The international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.

Bauer:1979:STR

[BBP^{+}79] F. L. Bauer, M. Broy, H. Partsch, P. Pepper, and H. Wössner. Systematics of transformation rules. In Bauer and Broy [BB79], pages 273–289. CODEN LNCSD9. ISBN 0-387-09251-X (print), 3-540-35312-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0069.htm;   http://www.springerlink.com/content/978-
0-387-09251-5;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=69;   https://
link.springer.com/book/10.1007/BFb0014649. The interna-

tional summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.

Bauer:1958:PUL

[BBR^{+}58] F. L. Bauer, H. Bottenbruch, H. Rutishauser, K. Samelson, P. Graeff, P. Lauchli, and M. Paul. Proposal for a universal language for the description of computing processes. In J. W. Carr, editor, Computer Programming and Artificial Intelligence, University of Michigan Summer School 1958, pages 355–373. ????, ????, 1958.

Bauer:1993:LAS

[BBS93] Friedrich L. Bauer, Wilfried Brauer, and Helmut Schwichtenberg, editors. Logic and algebra of specification. Proceedings of the NATO Advanced Study Institute, held in Marktoberdorf, Germany, July 23–August 4, 1991. NATO ASI Series. Series F. Computer and Systems Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-642-63448-6 (print), 3-642-58041-6 (e-book). vii + 442 pp. LCCN QA76.9.L63.

Bauer:1964:DNG

[BD64] F. L. Bauer and H.-J. Dreyer. Deutsche Normen. (German) [German standards]. Blätter der DGVFM, 6(4):556–564, April 1964. CODEN ???? ISSN 0012-0200.

Broy:2001:PTC

[BD01] Manfred Broy and Ernst Denert, editors. Pioneers and Their Contributions to Software Engineering: sd&m Conference on Software Pioneers, Bonn, June 28/29, 2001, Original Historic Contributions. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

Bauer:1974:CCA

[BDG^{+}74] F. L. Bauer, F. L. De Remer, M. Griffiths, U. Hill, J. J. Horning, C. H. A. Koster, W. M. McKeeman, P. C. Poole, and W. M. Waite. Compiler Construction: an Advanced Course, volume 21 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 3-540-06958-5 (print), 3-662-21549-7 (e-book). xiv + 621

pp. LCCN QA76.6 .C6281. URL https://link.springer.com/
book/10.1007/978-3-662-21549-4.

Bauer:1969:AEP

[BDS69] F. L. Bauer, E. Deutsch, and J. Stoer. Abschätzungen für die Eigenwerte positiver linearer Operatoren. (German) [Estimates for the eigenvalues of positive linear operators]. Linear Algebra and its Applications, 2(3):275–301, July 1969. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0024379569900317.

Bauer:1975:SEA

[BDW^{+}75] F. L. Bauer, J. B. Dennis, W. M. Waite, C. C. Gotlieb, R. M. Graham, M. Griffiths, H. J. Helms, B. Morton, P. C. Poole, and D. Tsichritzis, editors. Software Engineering: An Advanced Course, volume 30 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. ISBN 0-387-08364-2, 3-540-07168-7, 3-540-08364-2. xii + 545 pp. LCCN QA76.6 .A33 1972.

Bauer:1976:CCA

[BE76] Friedrich Ludwig Bauer and J. Eickel. Compiler construction: an advanced course, volume 21 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1976. CODEN LNCSD9. ISBN 0-387-08046-5 (USA), 3-540-08046-5 (Berlin), 0-387-07542-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 638 pp. LCCN QA76.6 .C628 1976. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0021.htm;   http://www.springerlink.com/content/978-
0-387-08046-8;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=21.

Bauer:1987:MPC

[BEH^{+}87] Friedrich Ludwig Bauer, H. Ehler, A. Horsch, B. Möller, H. Partsch, O. Paukner, and P. Pepper, editors. The Munich Project CIP. Volume II: The program transformation system CIP-S, volume 292 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-18779-0 (New York), 3-540-18779-0 (Berlin). viii + 522 pp. LCCN QA76.6 .M86 1985.

Bauer:1981:KS

[BEPP81] Friedrich L. Bauer, Andrei P. Ershov, M. Paul, and Alan J. Perlis. Klaus Samelson. Acta Informatica, 15(1):1–2, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:1958:NFP

[BF58] Friedrich L. Bauer and Evelyn Frank. Note on formal properties of certain continued fractions. Proceedings of the American Mathematical Society, 9:340–347, 1958. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL https://

www.jstor.org/stable/2032983.

Bauer:1960:NET

[BF60] F. L. Bauer and C. T. Fike. Norms and exclusion theorems. Numerische Mathematik, 2:137–141, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baumann:1964:IA

[BFBS64] R. Baumann, M. Feliciano, F. L. Bauer, and K. Samelson. Introduction to Algol. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1964. x + 142 pp. LCCN QA76.5 .B33.

Bauer:1951:BST

[BG51] F. L. Bauer and H. Görtler. Bücher: Sommerfeld: Theoretische Physik/Frössel: Kompressible Strömung an gewölbter Wand/ Müller: Ergänzungen und Berichtigungen. Physik Journal, 7(10): 477–479, October 1951. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Bauer:1971:IEUa

[BG71a] Friedrich L. Bauer and Gerhard Goos. Informatik: Eine einführende Übersicht, Teil I, volume 80 of Heidelberger Taschenbücher (Sammlung Informatik). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05303-4, 3-540-05303-4. xii + 213 pp. LCCN QA76 .B319.

Bauer:1971:IEUb

[BG71b] Friedrich L. Bauer and Gerhard Goos. Informatik: Eine einführende Übersicht, Teil II, volume 91 of Heidelberger Taschenbücher (Sammlung Informatik). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05487-1, 3-540-05487-1. xii + 202 pp. LCCN QA76 .B319.

Bauer:1973:IT

[BG73] Friedrich L. Bauer and Gerhard Goos. Informatik: Eine einführende Übersicht. Erster Teil. (German) [Computer Science: An Introductory Review. First part]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. ISBN 3-642-96161-4. xii + 220 pp. LCCN QA75.5-

76.95. URL https://link.springer.com/book/10.1007/978-

3-642-96161-8. Eine einführende Übersicht, Zweite Auflage, Heidelberger Taschenbücher, Band 80. (Sammlung Informatik).

Bauer:1974:IEU

[BG74] Friedrich L. Bauer and Gerhard Goos. Informatik: Eine einführende Übersicht. Zweiter Teil. (German) [Computer Science: an Introductory Review. Second part]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 0-387-06899-6, 3-540-06899-6. xiii + 207 pp. LCCN

QA76 .B3192. URL https://link.springer.com/book/10.

1007/978-3-642-96224-0. Eine einführende Übersicht, Zweite Auflage, Heidelberger Taschenbücher, Band 91. (Sammlung Informatik).

Bauer:1976:CIC

[BG76] F. L. Bauèr and G. Gooz. Informatika: vvodnyĭ kurs. Izdat. “Mir”, Moscow, USSR, 1976. 484 pp. Translated from the German by V. K. Sabel′feld, Edited by A. P. Eršov.

Bauer:1982:IET

[BG82] Friedrich L. Bauer and Gerhard Goos. Informatik. Erster Teil. (German) [Computer science. First part], volume 80 of Heidelberger Taschenbücher [Heidelberg Paperbacks]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1982. ISBN 3-540-11722-9. xviii + 320 pp. LCCN QA76.6-76.66. Eine einführende Übersicht. [An introductory survey].

Bauer:1984:IEU

[BG84] Friedrich L. Bauer and Gerhard Goos. Informatik. Eine einführende Übersicht. Zweiter Teil. (German) [Computer science. An introductory overview. Second part], volume 91 of Heidelberger Taschenbücher. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13121-3 (print), 3-642-96811-2 (e-book). xvi + 345 + 135 + 4 pp. LCCN QA76.6-76.66.

Bauer:1990:CIC

[BG90] F. L. Bauèr and G. Gooz. Informatika. Chast′ 1, 2. “Mir”, Moscow, Russia, second edition, 1990. ISBN 5-03-000298-7. Part 1: 324 + x pp.; Part 2: pp. 325–744 pp. Vvodnyĭ kurs. [An introductory survey], Translated from the German by M. K. Valiev, V. G. Kerbel′and V. K. Sabel′fel′d, Translation edited by A. P. Ershov, With a preface by Ershov and Sabel′fel′d.

Bauer:1991:IEU

[BG91] Friedrich L. Bauer and Gerhard Goos. Informatik 1. Ein einführende Übersicht. (German). (German) [Computer Science 1. An Introductory Overview]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth revised edition, ???? 1991. ISBN 3-540-52974-8. xviii + 393 pp. LCCN QA9 .B296 1991.

Bauer:1992:IEU

[BG92] Friedrich L. Bauer and Gerhard Goos. Informatik 2. Eine einführende Übersicht. (German) [Computer Science 2. An Introductory Overview]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1992. ISBN 0-387-55567-6, 3-540-55567-6 (print), 3-642-58154-4. xvi + 345 pp. LCCN QA75.5-76.95.

Bauer:1975:IAL

[BGH75] Friedrich L. Bauer, Rupert Gnatz, and Ursula Hill. Informatik: Aufgaben und Lösungen zum Erster Teil. (German) [Computer science: exercises and solutions for part I], volume 159 of Heidelberger Taschenbücher (Sammlung Informatik). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. ISBN 3-540-07007-9 (print), 3-642-66005-3 (e-book). ISSN 0073-1684. xi + 163 pp. (loose errata)

pp. LCCN QA75.5-76.95. URL http://link.springer.com/
openurl?genre=book&isbn=978-3-540-07007-8.

Bauer:1976:IAL

[BGH76] Friedrich L. Bauer, Rupert Gnatz, and Ursula Hill. Informatik: Aufgaben und Lösungen. Teil II. (German) [Computer science: exercises and solutions. Part II], volume 160 of Heidelberger Taschenbücher (Sammlung Informatik). Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 3-642-66065-7. x + 173 pp. LCCN QA75.5-

76.95. URL http://www.springerlink.com/content/978-3-
642-66065-8.

Bauer:1981:ZZZ

[BGH81] Friedrich L. Bauer, Rupert Gnatz, and Ursula Hill. Zbiór zadań z informatyki. (Polish) [A collection of problems in information science]. Seria Informatyka. [Information Science Series]. Wydawnictwa Naukowo-Techniczne (WNT), Warszawa, Poland, 1981. ISBN 83-204-0250-6. 349 pp. Translated from the German by Jarosław Cichorski.

Bauer:1979:ES

[BGHS79] F. L. Bauer, G. H. Golub, A. S. Householder, and K. Samelson. Eduard L. Stiefel: 4/21/1909–11/27/1978. Numerische Mathematik, 32(4):480–481, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). With a German translation.

Bauer:1959:C

[BH59] F. L. Bauer and A. S. HouseHolder. Correspondence. The Computer Journal, 2(3):102, ???? 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/3/102.full.pdf+html.

Bauer:1960:MCR

[BH60a] F. L. Bauer and A. S. Householder. Moments and characteristic roots. Numerische Mathematik, 2(1):42–53, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1960:SII

[BH60b] F. L. Bauer and A. S. Householder. Some inequalities involving the Euclidean condition of a matrix. Numerische Mathematik, 2 (1):308–311, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1961:ANC

[BH61] F. L. Bauer and A. S. Householder. Absolute norms and characteristic roots. Numerische Mathematik, 3(1):241–246, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1987:MJH

[BH87] F. L. Bauer and A. S. Householder. In memoriam: J. H. Wilkinson (1919–1986). Numerische Mathematik, 51(1):1–2, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:2008:CFG

[BH08] Friedrich L. Bauer and Christoph Haenel. Carl Friedrich Gauß, das 17-Eck und MATHEMATICA. (German) [Carl Friedrich Gauß, the 17-gon and Mathematica]. Informatik Spektrum, 31(5):??, ???? 2008. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s00287-008-0268-7.

Bauer:1950:QEK

[BHB50] F. L. Bauer, J. V. Harlem, and W. Braunbek. Zur Quantentheorie der Elementarteilchen/Kernsplitterung bei Antimon/ Streuversuche mit energiereichsten Protonen und Neutronen. (German) [On the quantum theory of elementary particles/ nuclear fragmentation in antimony/scattering experiments with the most energetic protons and neutrons]. Physik Journal, 6(11): 513–514, November 1950. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Bauer:1967:DA

[BHO^{+}67] Friedrich L. Bauer, Alston S. Householder, Frank W. J. Olver, Heinz Rutishauser, Klaus Samelson, and Eduard Stiefel. Description of ALGOL 60, volume 1a of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. xii +

326 pp. LCCN QA76.5 .R87. URL http://www.springerlink.
com/content/978-3-642-86934-1.

Bauer:1956:BBF

[BHS^{+}56] F. L. Bauer, R. Hüper, G. Schumann, O. Rang, W. Hanle, J. van Calker, G. Kurtze, and W. Oppelt. Bücher Bücher: Friedrichs:

Quantum Theory of Fields/Kaplan: Nuclear Physics/Bechert, Gerthsen: Atomphysik/Sears, Zemansky: University Physics/ Mayer: Physik dünner Schichten/Raaz, Köhler: Bau und Bildung der Kristalle/Przibram: Verfärbung und Lumineszenz/ Smyth: Dielectric Behavior and Structure/Kudrjawzew: Anwendung von Ultraschallverfahren bei physikalischchemischen Untersuchungen/Davis, Weed: Grundlagen der Industriellen Elektronik. Physik Journal, 12(3):137–143, March 1956. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Bauer:1965:MRE

[BHSS65] Friedrich L. Bauer, Josef Heinhold, Klaus Samelson, and Robert Sauer. Moderne Rechenanlagen; eine Einführung. (German) [Modern computing systems; an introduction], volume 5 of Leitfäden der angewandten Mathematik und Mechanik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1965. 357 pp. LCCN QA76 .B32.

Belaghi:2013:BMT

[BKB13] Mahmoud Jafari Shah Belaghi, Sergey Khrushchev, and Agamirza E. Bashirov. On Bauer–Muir transform of continued fractions. International Journal of Number Theory (IJNT), 9(2):321–332, March 2013. ISSN 1793-0421 (print), 1793-7310

(electronic). URL https://www.worldscientific.com/doi/
10.1142/S1793042112501369.

Bauer:1989:FPC

[BMPP89] Friedrich Ludwig Bauer, Bernhard Möller, Helmut Partsch, and Peter Pepper. Formal program construction by transformationscomputer-aided, intuition-guided programming. IEEE Transactions on Software Engineering, 15(2):165–180, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=21743.

Bode:2004:FB

[Bod04] Arndt Bode. F. L. Bauer 80. Informatik Spektrum, 27(4):309, August 2004. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Bauer:1971:IPD

[BR71] F. L. Bauer and C. Reinsch. Inversion of positive definite matrices by the Gauss–Jordan method. In J. H. Wilkinson and C. Reinsch, editors, Linear Algebra, volume II of Handbook for Automatic Computation, pages 45–49. Springer-Verlag,

Berlin, 1971. URL https://link.springer.com/chapter/
10.1007/978-3-642-86940-2_3;   https://link.springer.
com/chapter/10.1007/978-3-662-39778-7_3. Contribution

I/3.

Broy:2015:LDI

[Bro15] Manfred Broy. A life dedicated to informatics: an obituary for Prof. Friedrich L. Bauer. Acta Informatica, 52(4– 5):299–301, June 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/article/10.1007/s00236-015-0241-2;   http://link.
springer.com/content/pdf/10.1007/s00236-015-0241-2.pdf.

Bauer:1963:NAN

[BRS63] F. L. Bauer, H. Rutishauser, and E. Stiefel. New aspects in numerical quadrature. In Proceedings of the Symposium on Applied Mathematics, Vol. XV, pages 199–218. American Mathematical Society, Providence, RI, USA, 1963.

Bruderer:2011:KZE

[Bru11] Herbert Bruderer. Konrad Zuse und die ETH Zürich. (German) [Konrad Zuse and the ETH Zurich]. Informatik Spektrum, 34(6): ??, ???? 2011. CODEN INSKDW. ISSN 0170-6012 (print), 1432-

122X (electronic). URL http://link.springer.com/article/
10.1007/s00287-011-0573-4.

Bauer:1957:VAV

[BS57a] F. L. Bauer and K. Samelson. Verfahren zur automatischen Verarbeitung von kodierten Daten und Rechenmaschine zur Ausübung des Verfahrens. (German) [Method for the automatic processing of coded data and calculating machine to carry out the method]. Report B44 122 IX/42m, ????, ????, 1957. Auslegeschrift Nr. 1094019 vom 1.12.1960.

Bauer:1957:PIG

[BS57b] Friedrich L. Bauer and Klaus Samelson. Polynomkerne und Iterationsverfahren. (German) [Polynomial kernels and iteration methods]. Mathematische Zeitschrift, 67:93–98, 1957. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic).

Bauer:1959:NFT

[BS59] F. L. Bauer and H. J. Stetter. Zur numerischen Fourier-Transformation. (German) [On the numerical Fourier transformation]. Numerische Mathematik, 1(1):208–220, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0029-599X&volume=1&issue=1&spage=
208.

Bauer:1960:PCL

[BS60a] F. L. Bauer and K. Samelson. The problem of a common language, especially for scientific numerical work (motives, restrictions, aims and results of the Zurich Conference on ALGOL). In Anonymous, editor, Information Processing 1959: Proceedings of the International Conference on Information Processing, UNESCO, Paris, 15–20 June 1959, pages 120–125. UNESCO, Paris, France, 1960.

Bauer:1960:VAV

[BS60b] Friedrich Ludwig Bauer and Klaus Samelson. Verfahren zur automatischen Verarbeitung von kodierten Daten und Rechenmaschine zur Ausübung des Verfahrens. (German) [Method for the automatic processing of coded data and calculating machine to carry out the method]. German patent 1,094,019., December

1, 1960. URL https://link.springer.com/chapter/10.1007/
978-3-642-48354-7_2;   https://patents.google.com/patent/

DE1094019B. Patent filed 30 March 1957.

Bauer:1962:ACM

[BS62a] F. L. Bauer and K. Samelson. Automatic computing machines and method of operation. US Patent 3,047,228., July 31, 1962. US Patent Application filed 28 March 1958.

Bauer:1962:MVP

[BS62b] F. L. Bauer and K. Samelson. Maschinelle Verarbeitung von Programmiersprachen. (German) [Machine processing of programming languages]. In Hoffmann [Hof62], pages 227–268. ISBN 3-322-96126-5 (print), 3-322-96260-1. LCCN QA75.5-

76.95. URL https://link.springer.com/book/10.1007/978-
3-322-96260-7.

Bauer:1962:ANM

[BS62c] F. L. Bauer and J. Stoer. Algorithm 105: Newton Maehly. Communications of the ACM, 5(7):387–388, July 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1976:LHI

[BS76] Friedrich Ludwig Bauer and Klaus Samelson, editors. Language hierarchies and interfaces: international summer school [held July 23 to August 2, 1975 in Marktoberdorf], volume 46 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L335. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0046.htm;   http://www.springerlink.com/content/978-
0-387-07994-3;   http://www.springerlink.com/openurl.
asp?genre=book&isbn=978-3-540-07994-1;   http://www.
springerlink.com/openurl.asp?genre=issue&issn=0302-9743&
volume=46.

Bauer:1981:ERS

[BS81] F. L. Bauer and G. Schmidt. Erinnerungen an Robert Sauer. Beiträge zum Gedächtniskolloquium anläßlich seines 10. Todestages. (German) [Memories of Robert Sauer. Contributions to the Memorial Colloquium on the occasion of the 10th anniversary of his death]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10951-X, 3-642-93190-1. 75 pp. LCCN QA75.5-76.95.

Bauer:2000:FSC

[BS00] Friedrich L. Bauer and Ralf Steinbruggen, editors. Foundations of secure computation. 20th international summer school, Marktoberdorf, Germany, July 27–August 8, 1999, volume 175 of NATO ASI Series. Series F. Computer and Systems Sciences. IOS Press and Omsha, Amsterdam, The Netherlands and Tokyo, Japan, 2000. ISBN 1-58603-015-9, 4-274-90350-8. LCCN QA76.9.A25 F66 2000.

Bauer:1961:AMN

[BSW61] F. L. Bauer, J. Stoer, and C. Witzgall. Absolute and monotonic norms. Numerische Mathematik, 3:257–264, December 1961.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1972:ZPV

[BW72a] F. L. Bauer and H. Wössner. Zuses “Plankalkül”, ein Vorläufer der Programmiersprachen — gesehen vom Jahre 1972. (German) [Zuses “Plankalkül”, a forerunner of programming languages, seen from the year 1972]. Elektronische Rechenanlagen, 14(??): 111–118, ???? 1972. CODEN ELRAA4. ISSN 0013-5720.

Bauer:1972:PKZ

[BW72b] Friedrich L. Bauer and H. Wössner. The “Plankalkül” of Konrad Zuse: a forerunner of today’s programming languages. Communications of the ACM, 15(7):678–685, July 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Bauer:1981:ASP

[BW81] F. L. Bauer and H. Wössner. Algorithmische Sprache und Programmentwicklung. Unter Mitarbeit von H. Partsch und P. Pepper. (German) [Algorithmic language and program development. With the collaboration of H. Partsch and P. Pepper]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-09853-4, 3-540-09853-4. xv + 513 pp. LCCN QA76.6 .B395 1981.

Bauer:1982:ALP

[BW82] F. L. Bauer and H. Wössner. Algorithmic Language and Program Development. Texts and Monographs in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-11148-4, 3-540-11148-4. xvi + 497 pp. LCCN QA76.6 .B39513 1982. In memorian Klaus Samelson (1918–1980). In collaboration with H. Partsch and P. Pepper. Translated from the German original [BW81].

Bauer:1988:CEA

[BW88] Friedrich L. Bauer and Martin Wirsing. Crypt-equivalent algebraic specifications. Acta Informatica, 25(2):111–153, February 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:1991:EAG

[BW91] Friedrich L. Bauer and Martin Wirsing. Elementare Aussagenlogik. (German) [Elementary propositional logic]. Mathematik für Informatiker. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-52974-8 (print), 3-642-84263-1 (e-book). x + 228 pp. LCCN QA9 .B296 1991.

Carr:1966:BRI

[Car66] John W. Carr III. Book review: Introduction to ALGOL (R. Baumann, M. Feliciano, F. L. Bauer and K. Samelson). SIAM Review, 8(3):406–407, July 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://www.jstor.

org/stable/2028240.

Calmet:1996:AIS

[CCP96] Jacques Calmet, J. A. Campbell, and Jochen Pfalzgraf, editors. Artificial intelligence and symbolic mathematical computation: international conference, AISMC-3, Steyr, Austria, September 23–25, 1996, proceedings, volume 1138 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-61732-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .A766 1996.

URL http://link.springer-ny.com/link/service/series/
0558/tocs/t1138.htm;   http://link.springer.com/book/
10.1007/3-540-61732-9;   http://www.springerlink.com/
content/978-3-540-61732-7;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=1138.

Third International Conference on Artificial Intelligence and Symbolic Mathematical Computation (AISMC-3), held in Steyr (Austria), September 23–25, 1996.

Chen:1972:SSE

[Che72] Etang Chen. Structure of simplexes of equilibrium states in quantum statistical mechanics. Journal of Mathematical Physics, 13 (8):1130–1135, August 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v13/i8/p1130_s1.

Chen:1974:CSS

[Che74] Etang Chen. Choquet simplexes of states in physical systems. Journal of Mathematical Physics, 15(11):2022–2025, November 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v15/i11/p2022_s1.

Christensen:2007:RDS

[Chr07] Chris Christensen. Review of “Decrypted Secrets: Methods and Maxims of Cryptology”, Fourth Edition by F. L. Bauer. Cryptologia, 31(3):281–283, July 2007. CO-DEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (elec-

tronic). URL http://www.informaworld.com/smpp/content%
7Econtent=a779878432. See [Bau07d].

Church:1951:FBA

[Chu51] Alonzo Church. Review: F. L. Bauer. Zur Algebraik des Logikkalküls. Methodos, vol. 1 (1949), pp. 288–292. Journal of Symbolic Logic, 16(1):62, March 1951. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL https://www.

jstor.org/stable/2268681.

Culik:1972:BRJb

[Čul72a] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 4 (1962–1963), pp. 420–453; also ibid., vol. 6 (1963), pp. 1–17; and The Computer Journal, vol. 5 no. 4 (1963), pp. 349–367. Journal of Symbolic Logic, 37(3): 623–624, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL https://www.jstor.org/

stable/2272794.

Culik:1972:BRJa

[Čul72b] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur (editor), A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 2 (1960), pp. 106–136; also Communications of the Association for Computing Machinery, vol. 3 (1960), pp. 299–314. Journal of Symbolic Logic, 37(3):622–623, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL ttps://www.jstor.

org/stable/2272793.

Dijkstra:1971:RBG

[Dij71] Edsger W. Dijkstra. [review of] Bauer, Goos: “Informatik, Erster Teil” Heidelberger Taschenbücher, Band 80, Springer Verlag, 1971. Circulated privately., May 1971. URL http://www.cs.

utexas.edu/users/EWD/ewd03xx/EWD311.PDF.

Dijkstra:1972:BRB

[Dij72] Edsger W. Dijkstra. [book review of: Bauer–Goos, Informatik, Zweiter Teil]. Circulated privately., March 1972. URL http://

www.cs.utexas.edu/users/EWD/ewd03xx/EWD330.PDF.

Dijkstra:1986:FBC

[Dij86] Edsger W. Dijkstra. F. L. Bauer’s Conjecture is F. L. Bauer’s Theorem. Circulated privately., August 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD970.PDF.

Dijkstra:1988:ZJI

[Dij88] Edsger W. Dijkstra. Zwanzig Jahre Institut für Informatik. (German) [Twenty years of the Institute for Informatics]: A new science, from birth to maturity. Report, ETH Zürich, Zürich, Switzerland, October 16, 1988. 25–38 pp. URL http:

//pascal.hansotten.com/uploads/wirth/Zwanzig%20Jahre%
20Institut%20Informatik%20095.pdf.

deLeeuw:2007:HIS

[dLB07] Karl de Leeuw and J. A. Bergstra, editors. The History of Information Security: a Comprehensive Handbook. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-51608-5 (hardcover). xi + 887 pp. LCCN Z103 .H63 2007. URL http://

www.sciencedirect.com/science/book/9780444516084.

Dijkstra:1976:CAP

[DS76] Edsger W. Dijkstra and C. S. Scholten. Comments on Arbeitsblatt 3 from o. Prof. Dr. F. L. Bauer. Circulated privately.,

1976. URL http://www.cs.utexas.edu/users/EWD/ewd06xx/
EWD602.PDF.

Deutsch:1975:BGG

[DZ75] E. Deutsch and C. Zenger. On Bauer’s generalized Gershgorin discs. Numerische Mathematik, 24(1):63–70, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ershov:1981:AMM

[EK81] A. P. Ershov and Donald E. Knuth, editors. Algorithms in modern mathematics and computer science: proceedings, Urgench, Uzbek SSR, September 16–22, 1979, volume 122 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CO-DEN LNCSD9. ISBN 0-387-11157-3 (New York), 3-540-11157-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA9.58 .A43 1981. URL https://link.springer.com/book/

10.1007/3-540-11157-3. The symposium was organized by the Academy of Sciences of the Uzbek S.S.R.

Eickel:1963:SCG

[EPBS63] J. Eickel, M. Paul, F. L. Bauer, and K. Samelson. A syntax controlled generator of formal language processors. Communications of the ACM, 6(8):451–455, August 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fassbender:2006:CVF

[FGH^{+}06] Heike Fassbender, Michael Griebel, Olga Holtz, G. W. (Pete) Stewart, and Christoph Zenger. Curriculum vitae of Friedrich Ludwig Bauer. Linear Algebra and its Applications, 417(2–3): 299–300, September 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special Issue in honor of Friedrich Ludwig Bauer.

Freiman:1972:IPP

[FGR72] C. V. Freiman, J. E. Griffith, and J. L. Rosenfeld, editors. Information processing 71: proceedings of IFIP Congress 71, Ljubjana, Yugoslavia, August 23–28, 1971, volume 1 of IFIP congress series. North-Holland, Amsterdam, The Netherlands, 1972. ISBN 0-7204-2063-6. LCCN QA75.5 .I532 1971. URL

https://dblp.org/db/conf/ifip/ifip71-1.html.

Feijen:1990:BOB

[FvGGM90] W. H. J. Feijen, A. J. M. van Gasteren, David Gries, and J. Misra, editors. Beauty is our Business: a Birthday Salute to Edsger W. Dijkstra. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. xix + 453 pp. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/en/

search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

G:1970:BRR

[G.70] W. G. Book review: R. Sauer, I. Szabo, T. P. Angelitch, G. Aumann, F. L. Bauer, R. Bulirsch, H. P. Kunzi, H. Rutishauser, K. Samelson, R. Sauer, and J. Stoer, Mathematische Hilfsmittel des Ingenieurs. III, Springer-Verlag, Berlin, 1969, xviv + 534 pp., 24 cm. Price $24.50. Mathematics of Computation, 24(110): 475, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL https://www.jstor.org/stable/
2004494.

Gander:2015:LCP

[Gan15] Walter Gander. The life of a computer pioneer [Friedrich L. Bauer]. Blogg@CACM Web site., April 13, 2015. URL http:

//cacm.acm.org/blogs/blog-cacm/185577-the-life-of-a-
computer-pioneer/fulltext.

Gotze:1971:RSG

[GB71] H. Götze and F. L. Bauer. Robert Sauer, geb. 16.9.1898 – gest. 22.8.1970. (German) Robert Sauer, born 16.9.1898 – died 22.8.1970. Numerische Mathematik, 16(4):i–ii, January 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gotze:1981:KSG

[GB81] H. Götze and F. L. Bauer. Klaus Samelson: geb. 21.12.1918, gest. 25.5.1980 (German) [Klaus Samelson: born 21.12.1981, died 25.5.1980]. Numerische Mathematik, 36(2):109, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gotze:1980:KSG

[GB81] H. Götze and F. L. Bauer. Klaus Samelson: geb. 21. 12. 1918, gest. 25. 5. 1980. Numerische Mathematik, 36(2):109, 1980/81. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Geppert:1966:BBJ

[Gep66] M. G. Geppert. Buchbesprechung: Backus, J. W., F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H.

Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. Van Wijngarden und M. Woodger: Bericht über die algorithmische Sprache Algol 60. (Report on the algorithmic language Algol 60. Übers. Kerner J. O. u. Mitarb.). Akademie-Verlag Berlin 1962. vi + 49 S., Preis MDN 10,50. Biometrische Zeitschrift, 8(1-2): 121, 1966. CODEN BIZEB3. ISSN 0006-3452.

Gheorghe:1987:BRM

[Ghe87] M. Gheorghe. Book review: The Munich Project CIP Volume I: The Wide Spectrum Language CIP L By the CIP Language Group. L.N.C.S. vol. 183 by F. L. Bauer, R. Berghammer, M. Broy, W. Dosch, F. Geiselbrechtinger, R. Gnatz, H. Hangel, W. Hesse, B. Krige-Brückner, A. Laut, T. Matzner, B. Möller, F. Nickl, H. Partsch, P. Pepper, K. Samelson, M. Wirsing, H. Wössner. Bulletin mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumanie Nouvelle Série, 31 (79)(4):378, 1987. ISSN 0007-4691. URL

https://www.jstor.org/stable/43681817.

Ghilardi:1974:BRA

[Ghi74] Gilberto Ghilardi. Book review: Advanced course in software engineering by F. L. Bauer. Giornale degli Economisti e Annali di Economia Nuova Serie, 33(11–12):891, November/December

1974. ISSN 0017-0097. URL https://www.jstor.org/stable/
23243304.

Ghizzetti:2010:AQA

[Ghi10] Aldo Ghizzetti, editor. Alcune questioni di analisi numerica. Lectures given at a Summer School of the Centro Internazionale Matematico Estivo (C.I.M.E.), held in Perugia, Italy, September 7–16, 1964. (Italian) [Some numerical analysis issues], volume 35 of C.I.M.E. Summer Schools. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN AINFA2. ISBN 3-642-11026-6 (print), 3-642-11027-4 (e-book). ISSN 0001-5903 (print), 1432-0525 (electronic).

LCCN QA297 .C56 2010. URL https://link.springer.com/
10.1007/978-3-642-11027-6.

Grau:1967:TA

[GHL67] A. A. Grau, U. Hill, and H. Langmaack. Translation of ALGOL 60, volume 137 of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-642-86937-8 (e-book), 3-642-86939-4 (print). ISSN 0072-7830 (print), 2196-9701 (electronic). ix +

403 pp. LCCN QA1-939. URL http://www.springerlink.com/
content/978-3-642-86937-2.

Gries:1990:BDO

[Gri90] David Gries. Binary to decimal, one more time. In Feijen et al. [FvGGM90], chapter 16, pages 141–148. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/en/

search/?an=0718.68004. This paper presents an alternate proof of Knuth’s algorithm [Knu90] for conversion between decimal and fixed-point binary numbers.

Gericke:1953:BBS

[GTB53] H. Gericke, G. L. Tautz, and F. L. Bauer. Bücher Bücher: Stabler: Mathematical Thought/Petrovskij: Integralgleichungen/ Peters: Sechsstellige Tafel der trigonometrischen Funktionen/ Sommerfeld: Thermodynamik und Statistik/Braun: Das Marsprojekt/Neher: Menschen zwischen den Planeten. Physik Journal, 9(12):569–573, December 1953. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Householder:1959:CME

[HB59] Alston S. Householder and Friedrich L. Bauer. On certain methods for expanding the characteristic polynomial. Numerische Mathematik, 1:29–37, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1960:CIM

[HB60] A. S. Householder and F. L. Bauer. On certain iterative methods for solving linear systems. Numerische Mathematik, 2 (1):55–59, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://www.osti.

gov/biblio/4158095-certain-iterative-methods-solving-
linear-systems.

Heinhold:1962:FPW

[HB62] J. Heinhold and F. L. Bauer. Fachbegriffe der Programmierungstechnik. Wörterverzeichnis für die Programmierung von Digital-Rechenanlagen mit Stichworten in fünf Sprachen. 2. verbess. und durch ein ALGOL-Wörterbuch erweiterte Auflage. (German) [Technical terms of programming technology. List of words for programming digital computers with key words in five languages. Second improved and extended edition with an ALGOL dictionary]. R. Oldenbourg Verlag, München, West Germany, 1962. 52 pp. LCCN QA76.5 .G4 1962.

Heinisch:1968:BKP

[Hei68] O. Heinisch. Buchbesprechung: Karlson, P.: Zauber der Zahlen. Eine unterhaltsame Mathematik für jedermann. Neu bearbeitet von F. L. Bauer. Verlag Ullstein, Darmstadt 1965. 432 S., zahlr. Abb., Leinen, DM 16,80. Biometrische Zeitschrift, 10(2):152, 1968. CODEN BIZEB3. ISSN 0006-3452.

Hoffmann:1962:DIP

[Hof62] Walter Hoffmann, editor. Digitale Informationswandler: Probleme der Informationsverarbeitung in ausgewählten Beiträgen. (German) [Digital information processors: selected articles on problems of information processing]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1962. ISBN 3-322-96126-5 (print), 3-322-96260-1. LCCN QA75.5-76.95. URL https://link.

springer.com/book/10.1007/978-3-322-96260-7.

Hohn:1964:BRI

[Hoh64] Franz E. Hohn. Book review: Introduction to ALGOL by R. Baumann, M. Feliciano, F. L. Bauer, K. Samelson. Pi Mu Epsilon Journal, 4(1):31, Fall 1964. CODEN PMEJBR. ISSN 0031-952X.

URL https://www.jstor.org/stable/24339270.

Ionescu:1984:BRP

[Ion84] Bueur B. Ionescu. Book review: A Programming approach to computability, Texts and Monographs in Computer Science by A. J. Kfoury, Robert N. Moll, Michael A. Arbib, David Gries, F. L. Bauer, K. S. Fu, J. J. Horning, R. Reddy, V. C. Tsichritzis, W. M. Waite. Bulletin mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumanie, Nouvelle Série, 28 (76)(2):188–189, ???? 1984. ISSN 0007-4691.

URL https://www.jstor.org/stable/43680898.

Kadner:1976:BFB

[Kad76] H. Kadner. Buchbesprechung: F. L. Bauer/G. Goos, Informatik, eine einführende Übersicht, Tl. II, 208 S., Berlin-Heidelberg-New York 1974. Springer Verlag. DM 14,80. Zeitschrift für Angewandte Mathematik und Mechanik, 56(1):63, 1976. CODEN ZA-MMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Kirschstein:1953:AAM

[KB53] G. Kirschstein and F. L. Bauer. Arbeitstagung des Ausschusses für Mechanisierung der Dokumentation in Berlin/Theoretischphysikalische Arbeitstagung in Marquartstein. (German) [Working conference of the Committee for Mechanization of Documentation in Berlin/Theoretical-physical working conference in Marquartstein]. Physik Journal, 9(2):84, February 1953. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kilov:2012:ROF

[Kil12] Haim Kilov. Review of Origins and Foundations of Computing, by Friedrich L. Bauer. ACM SIGACT News, 43(4):38–40, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Knodel:1966:BRF

[Knö66] W. Knödel. Book review: F. L. Bauer, J. Heinhold, K. Samelson, und R. Sauer, Moderne Rechenanlagen (Leitfäden der angewandten Mathematik und Mechanik: Band 5). Computing, 1(1): 88, March 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Knuth:1990:SPW

[Knu90] Donald E. Knuth. A simple program whose proof isn’t. In Feijen et al. [FvGGM90], chapter 27, pages 233–242. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/en/

search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Koblitz:2010:BRB

[Kob10] Neal Koblitz. Book review: Decrypted Secrets: Methods and Maxims of Cryptology. Fourth Edition. By Friedrich L. Bauer. SIAM Review, 52(4):777–779, December 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://www.jstor.org/stable/41062027.

Langer:1959:NAP

[Lan59] R. E. Langer, editor. On numerical approximation. Proceedings of a Symposium, Madison, April 21–23, 1958. The University of Wisconsin Press, Madison, WI, USA, 1959. x + 462 pp. LCCN QA3 .U45 no. 1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Langmaack:1969:LKG

[LB69] H. Langmaack and F. L. Bauer. Logik und Kybernetik. (German) [Logic and cybernetics]. Studium Generale, 22(??):137–150, ???? 1969. CODEN STGEAZ. ISSN 0039-4149.

Lehmann:1987:BRZ

[Leh87] N. J. Lehmann. Book review: Zuse, K., Der Computer — Mein Lebenswerk. Mit Geleitworten von F. L. Bauer und H. Zemanek. Berlin-Heidelberg-New York-Tokyo, Springer-Verlag 1984. xv + 218 S., 78 Abb., DM 39,–. ISBN 3-540-13814-5. Zeitschrift für Angewandte Mathematik und Mechanik, 67(3):211, 1987. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Levrie:1989:IMC

[Lev89] Paul Levrie. Improving a method for computing nondominant solutions of certain second-order recurrence relations of Poincarétype. Numerische Mathematik, 56(5):501–512, November 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1988:CBF

[Li88] Ren Cang Li. A converse to the Bauer–Fike type theorem. Linear Algebra and its Applications, 109(??):167–178, October 1988. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002437958890208X. This work refers to [BF60].

Livowschi:1981:BRP

[Liv81] L. Livowschi. Book review: Program Construction, Lect. Notes in Computer Science, 69 by F. L. Bauer, M. Broy. Bulletin mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumanie Nouvelle Série, 25 (73)(1): 104–106, 1981. ISSN 0007-4691. URL https://www.jstor.org/

stable/43680653.

Leap:2016:FIB

[LMNS16] Tom Leap, Tim McDevitt, Kayla Novak, and Nicolette Siermine. Further improvements to the Bauer–Millward attack on the Hill cipher. Cryptologia, 40(5):452–??, 2016. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Lipschutz-Yevick:1998:LE

[LYB98] Miriam Lipschütz-Yevick and F. L. Bauer. Letters to the editor. The Mathematical Intelligencer, 20(1):12–13, March 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.
1007/BF03024392;   https://link.springer.com/content/
pdf/10.1007/BF03024392.pdf.

Moler:1996:BRH

[MB96] Cleve Moler and Friedrich L. Bauer. Bauer remembers Householder and the Gatlinburg meetings. NA Digest, 96(27), July 18,

1996. URL http://www.maths.manchester.ac.uk/~higham/
conferences/householder/h96/bauertalk.pdf;   http://
www.netlib.org/na-digest-html/96/v96n27.html;   http:
//www3.math.tu-berlin.de/householder_2008/Cleve.htm. The

article contains notes for F. L. Bauer’s after-banquet talk entitled Memories of Alston Householder (1904–1993), Householder Symposium XIII, June 17–21, 1996 Pontresina, Switzerland.

Moritz:1956:BBM

[MBH^{+}56] E. Moritz, F. L. Bauer, A. Hinzpeter, H. Helbig, H. Wolf, W. Meyer-Eppler, A. Naumann, and H. Gartmann. Bücher Bücher: Morse, Feshbach: Methods of Theoretical Physics/ Feenberg, Pake: Angular Momentum/Zimmer: Quantenphysik und Atombau/Schulz-Methke: Photoelemente und Kristallphotozellen/Czech: Der Elektronenstrahl-Oszillograf/Reichardt: Grundlagen der Elektroakustik/Betz: Hydro und Aerodynamik/ Boyd, Seaton: Rocket Exploration of the Upper Atmosphere/Der Grosse Brockhaus. Physik Journal, 12(4):185–190, April 1956. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Moler:2015:CCF

[Mol15] Cleve Moler. Cleve’s Corner: Friedrich Bauer. MathWorks

Web site., June 1, 2015. URL https://blogs.mathworks.com/
cleve/2015/06/01/friedrich-bauer/.

Naur:1960:RALa

[NBB^{+}60] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Communications of the ACM, 3(5):299–314, May 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [AGG60]. Dedicated to the Memory of William Turanski [American group member who was killed by an automobile just prior to the January 1960 Conference].

Naur:1962:BAS

[NBB^{+}62] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, S. Ruti shauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Bericht über die algorithmische Sprache ALGOL 60. (German) [Report on the ALGOL 60 language]. Elektronisches Rechnen und Regeln. Sonderband 1. Akademie-Verlag, Berlin, Germany, 1962. xi + 49 pp. LCCN QA76.5 .N3515. Zur Erinnerung an William Turanski (In memory of William Turanski).

Naur:1963:RRA

[NBB^{+}63] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, and P. Nauer.

Revised report on the algorithmic language ALGOL 60. Communications of the ACM, 6(1):1–17, January 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Dedicated to the memory of William Turanski.

Naur:1965:RRA

[NBB^{+}65] P. Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL-60. (Russian), volume Moskau, Russia. Verlag “Mir”, 1965. 80 pp.

Naur:1997:RRA

[NBB^{+}97] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Kafz, J. Mc-Carthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language Algol 60. In Algol-like Languages, pages 19–49. Birkhäuser Boston Inc., Cambridge, MA, USA, 1997.

Nirschl:1964:BFV

[NS64] N. Nirschl and H. Schneider. The Bauer fields of values of a matrix. Numerische Mathematik, 6:355–365, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oquendo:2009:OTB

[OML09] W. F. Oquendo, J. D. Muñoz, and A. Lizcano. Oedometric test, Bauer’s law and the micro-macro connection for a dry sand. Computer Physics Communications, 180(4):616–620, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000046.

Pepper:1982:ADA

[PBB^{+}82] P. Pepper, M. Broy, F. L. Bauer, H. Partsch, W. Dosch, and M. Wirsing. Abstrakte Datentypen: Die algebraische Spezifikation von Rechenstrukturen. (German) [Abstract data types: the algebraic specification of arithmetic structures]. Informatik Spektrum, 5(??):107–119, ???? 1982. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Petzold:1999:FBH

[Pet99] Hartmut Petzold. F. L. Bauer als Historiker. (German) [F. L. Bauer as a historian]. Working document, Münchner Zentrum für Wissenschafts- und Technikgeschichte, Munich, Germany,

1999. URL https://webarchiv.typo3.tum.de/TUM/mzwtg/
fileadmin/w00bmt/www/Arbeitspapiere/petzold_bauer.pdf.

Pepper:1978:TAP

[PPWB78] P. Pepper, H. Partsch, H. Wössner, and F. L. Bauer. A transformational approach to programming. In Program transformations: Proceedings of the 3rd International Symposium on Programmming, Paris, 1978, pages 248–262. ????, ????, ???? 1978.

Pryde:1992:BFT

[Pry92] A. J. Pryde. A Bauer–Fike theorem for the joint spectrum of commuting matrices. Linear Algebra and its Applications, 173: 221–??, August 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). This work refers to [BF60].

Rutishauser:1955:DVP

[RB55] Heinz Rutishauser and Friedrich L. Bauer. Détermination des vecteurs propres d’une matrice par une méthode itérative avec convergence quadratique. (French) [Determination of the eigenvectors of a matrix by an iterative method with quadratic convergence]. Comptes rendus de l’Académie des sciences, Paris, 240: 1680–1681, 1955. ISSN 0001-4036 (print), 2419-6304 (electronic).

Reinsch:1968:RTN

[RB68] C. Reinsch and F. L. Bauer. Rational QR transformation with Newton shift for symmetric tridiagonal matrices. Numerische Mathematik, 11(3):264–272, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 257–272].

Reingold:1995:LE

[RBHM95] Edward M. Reingold, F. L. Bauer, Markus Heyerhoff, and Charles Musès. Letters to the editor. The Mathematical Intelligencer, 17(3):3–5, January 1995. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Rojas:2000:FCH

[RH00] Raúl Rojas and Ulf Hashagen, editors. The First Computers: History and Architectures. History of computing. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-18197-5 (hardcover), 0-585-35535-5 (electronic). xii + 457 pp. LCCN QA76.17 .F57 2000.

Rheinboldt:1965:IA

[Rhe65] Werner C. Rheinboldt. Book review: R. Baumann, M. Feliciano, F. L. Bauer, and K. Samelson, Introduction to ALGOL. Mathematics of Computation, 19(89):166, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2004138.

Rojas:1996:ARZ

[Roj96] Raúl Rojas. Die Architektur der Rechenmaschinen Z1 und Z3 von Konrad Zuse. (German) [The architecture of the calculating machines Z1 and Z3 by Konrad Zuse]. Informatik Spektrum, 19(6):??, ???? 1996. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s002870050041.

Rutishauser:1969:CAF

[Rut69] Heinz Rutishauser. Computational aspects of F. L. Bauer’s simultaneous iteration method. Numerische Mathematik, 13 (1):4–13, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://link.springer.

com/article/10.1007/BF02165269.

Rutishauser:2013:NAS

[Rut13] Hanna Rutishauser. Numerik, ALGOL und die Schweizer Hochalpen. Zur Arbeit an der Biografie von Heinz Rutishauser (1918–1970). (German) [Numerics, ALGOL and the Swiss High Alps. On the work on the biography of Heinz Rutishauser (1918–1970)]. Informatik Spektrum, 36(5):??, ???? 2013. CO-DEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00287-013-0730-z.

Samelson:1953:ORR

[SB53] Klaus Samelson and Friedrich L. Bauer. Optimale Rechengenauigkeit bei Rechenanlagen mit gleitendem Komma. (German) [Optimal calculation accuracy for calculators with a floating point]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 4(4):312–316, July 1953. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Samelson:1954:MNE

[SB54] Klaus Samelson and Friedrich L. Bauer. Maßnahmen zur Erzielung kurzer und übersichtlicher Programme für Rechenautomaten. (German) [Measures to achieve short and clear programs for calculators]. Zeitschrift für Angewandte Mathematik und Mechanik, 34(7):262–272, 1954. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Samelson:1959:SFG

[SB59] K. Samelson and F. L. Bauer. Sequentielle Formelübersetzung. (German) [Sequential formula translation]. Elektronische Rechenanlagen, 1(??):176–182, 1959. CODEN ELRAA4. ISSN 0013-5720. See also [SB60].

Samelson:1960:SFT

[SB60] K. Samelson and Friedrich L. Bauer. Sequential formula translation. Communications of the ACM, 3(2):76–83; addendum, 351, February 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Addendum, 351.

Samelson:1962:AP

[SB62] K. Samelson and F. L. Bauer. The ALCOR project. In ????, editor, Symbolic Languages in Data Processing. Proceedings of the ICC Symposium, Rome, March 26–31, 1962, pages 207–217. Gordon and Breach, New York, NY, USA, 1962.

Samelson:1983:SFT

[SB83] K. Samelson and Friedrich L. Bauer. Sequential formula translation (reprint). Communications of the ACM, 26(1):9–13, 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprint of [SB60].

Stoer:1990:NMG

[SB90] Josef Stoer and Roland Bulirsch. Numerische Mathematik. 2. (German) [Numerical mathematics. 2]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1990. ISBN 0-387-51482-1 (paperback), 3-540-51482-1 (paperback). xiv + 341 pp. LCCN QA297 .S823 1990. Eine Einführung—unter Berücksichtigung von Vorlesungen von F. L. Bauer. [An introduction, with reference to lectures by F. L. Bauer].

Samelson:2001:SFG

[SB01] K. Samelson and F. L. Bauer. Sequentielle Formelübersetzung. (German) [Sequential formula translation]. In Broy and Denert [BD01], pages 1–27. ISBN 3-540-42290-0 (print), 3-642-48354-2

(e-book). LCCN QA76.758. URL http://link.springer.com/
chapter/10.1007/978-3-642-48354-7_1/;   http://link.
springer.com/content/pdf/bfm:978-3-642-48354-7/1.pdf.

Siefkes:1999:FB

[SBE^{+}99a] Dirk Siefkes, Anette Braun, Peter Eulenhöfer, Heike Stach, and Klaus Städtler. Friedrich L. Bauer. In Pioniere der Informatik: Ihre Lebensgeschichte im Interview. (German) [Pioneers of computer science: their life stories in interviews] [SBE^{+}99b], chapter 3, pages 61–105. ISBN 3-642-58599-X, 3-642-63660-8 (print).

LCCN QA76.17. URL http://link.springer.com/openurl?
genre=book&isbn=978-3-642-63660-8.

Siefkes:1999:PII

[SBE^{+}99b] Dirk Siefkes, Anette Braun, Peter Eulenhöfer, Heike Stach, and Klaus Städtler. Pioniere der Informatik: Ihre Lebensgeschichte im Interview. (German) [Pioneers of computer science: their life stories in interviews]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-642-58599-X, 3-642-63660-8 (print). LCCN QA76.17. URL http:/

/link.springer.com/openurl?genre=book&isbn=978-3-642-
63660-8.

Smillie:1985:RFM

[SBET85] K. W. Smillie, F. L. Bauer, Ralph Erskine, and Henry S. Tropp. Reviews: O. I. Franksen, Mr. Babbage’s Secret; F. H. Hinsley, British Intelligence in the Second World War; T. M. Thompson, From Error-Correcting Codes Through Sphere Packings to Simple Groups; capsule reviews. Annals of the History of Computing, 7(2):185–191, April/June 1985. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1985/pdf/a2185.pdf;   http://www.computer.
org/annals/an1985/a2185abs.htm.

Schmid:1969:BRF

[Sch69] E. Schmid. Book review: F. L. Bauer and A. S. Householder and F. J. Olver and H. Rutishauser and K. Samelson and E. Stiefel, Handbook for Automatic computation, Vol. I Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete, Band 135. Computing, 4(3):276, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schmid:1970:BRF

[Sch70] E. Schmid. Book review: F. L. Bauer and A. S. Householder and F. W. J. Oliver and H. Rutishauser and K. Samelson and E. Stiefel (Hg.), Handbook for Automatic Computation, Vol. I, Part B (Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete: Band 137). Computing, 6(1–2):214, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Scrivener:1979:BRF

[Scr79] Steve Scrivener. Book review: F. L. Bauer, Software Engineering: An Advanced Course. Leonardo (Oxford, England), 12(2):171, 1979. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282

(electronic). URL ttps://www.jstor.org/stable/1573870.

Stoer:1993:NMG

[Sto93] Josef Stoer. Numerische Mathematik. 1. (German) [Numerical mathematics. 1]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., sixth edition, 1993. ISBN 0-387-56213-3, 3-540-56213-3. xiv + 314 pp. LCCN QA297-299.4. Eine Einführung— unter Berücksichtigung von Vorlesungen von F. L. Bauer. [An introduction—based on the lectures of F. L. Bauer].

Stoer:1994:NMG

[Sto94] Josef Stoer. Numerische Mathematik. 1. (German) [Numerical mathematics. 1]. Springer-Lehrbuch. [Springer Textbook]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., seventh edition, 1994. ISBN 0-387-57823-4, 3-540-57823-4, 3-662-09023-6. xii + 367 pp. LCCN QA297 .S78 1994. Eine Einführung—unter Berücksichtigung von Vorlesungen von F. L. Bauer. [An introduction—based on the lectures of F. L. Bauer].

Svecoa:1970:BSC

[Šve70] H. Švecoá. On the Bauer’s scaled condition number of matrices arising from approximate conformal mapping. Numerische Mathematik, 14(5):495–507, April 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:/

/www.springerlink.com/openurl.asp?genre=article&issn=
0029-599X&volume=14&issue=5&spage=495.

Shi:2012:SVB

[SW12] Xinghua Shi and Yimin Wei. A sharp version of Bauer–Fike’s theorem. Journal of Computational and Applied Mathematics, 236(13):3218–3227, July 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042712000787.

This work refers to [BF60].

vanWijngaarden:1969:RALb

[vWMPK69a] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Report of Algorithmic Language ALGOL 68. Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. ISBN 3-662-38647-X (print), 3-662-39504-5 (e-book). 79–218 pp. LCCN

QA76.9.A43. URL http://www.springerlink.com/content/
978-3-662-39504-2.

vanWijngaarden:1969:RALa

[vWMPK69b] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Report on the algorithmic language ALGOL 68. Numerische Mathematik, 14(2):79–218, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://www.fh-jena.de/~kleine/history/languages/
Algol68-Report.pdf;   http://www.fh-jena.de/~kleine/
history/languages/Algol68-RevisedReport.pdf.

Chu:2003:PEM

[wC03] Eric King wah Chu. Perturbation of eigenvalues for matrix polynomials via the Bauer–Fike theorems. SIAM Journal on Matrix Analysis and Applications, 25(2):551–573, January 2003. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). This work refers to [BF60].

Chu:2004:PEP

[wCL04] Eric King wah Chu and Wen-Wei Lin. Perturbation of eigenvalues for periodic matrix pairs via the Bauer–Fike theorems. Linear Algebra and its Applications, 378(1):183–202, February 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). This work refers to [BF60].

Chu:1986:GBF

[wEC86] King wah Eric Chu. Generalization of the Bauer–Fike Theorem. Numerische Mathematik, 49(6):685–691, November 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). This work refers to [BF60].

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1988/pdf/a4348.pdf;   http://www.computer.
org/annals/an1988/a4348abs.htm.

Wilkinson:1984:NMQ

[Wil84] J. H. Wilkinson. On neighbouring matrices with quadratic elementary divisors. Numerische Mathematik, 44(1):1–21, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). In honour of the 60th birthday of F. L. Bauer.

Wossner:1979:STT

[WPPB79] H. Wössner, P. Pepper, H. Partsch, and F. L. Bauer. Special transformation techniques. In Bauer and Broy [BB79], pages 290–321. CODEN LNCSD9. ISBN 0-387-09251-X (print), 3-540-35312-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL https://link.springer.

com/chapter/10.1007/BFb0014673. The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6, 3-662-38854-5, 3-662-39778-1. viii + 439 pp. LCCN QA251

.W67. URL http://www.springerlink.com/content/978-3-
662-39778-7.

Zenger:1968:CPB

[Zen68] Chr. Zenger. On convexity properties of the Bauer field of values of a matrix. Numerische Mathematik, 12(2):96–105, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).